### v2

# Semantic Web technologies in sensor-based personal health monitoring systems: A systematic mapping study

Abstract
In recent years, there has been an increased focus on early detection, prevention, and prediction of diseases. This, together with advances in sensor technology and the Internet of Things, has led to accelerated efforts in the development of personal health monitoring systems. This study analyses the state of the art in the use of Semantic Web technologies in sensor-based personal health monitoring systems. Using a systematic approach, a total of 43 systems are selected as representative of the current state of the art. We critically analyse the extent to which the selected systems address seven key challenges: interoperability, context awareness, situation detection, situation prediction, decision support, explainability, and uncertainty handling. We discuss the role and limitations of Semantic Web technologies in managing each challenge. We then conduct a quality assessment of the selected systems based on the data and devices used, system and components development, rigour of evaluation, and accessibility of research outputs. Finally, we propose a reference architecture to provide guidance for the design and development of new systems. This study provides a comprehensive mapping of the field, identifies inadequacies in the state of the art, and provides recommendations for future research.
Semantic Web technologies,
ontologies,
knowledge graphs,
linked data,
sensors,
Internet of Things,
health monitoring,
keywords:
volume: 0
1 Introduction
Non-communicable diseases are on the rise globally, resulting not only in decreased quality of life but also increasing healthcare costs [1]. For this reason, there have been accelerated efforts to develop personal health monitoring systems for early detection, prediction, and prevention of diseases. The emerging paradigm of precision health goes beyond treating existing diseases and rather focuses on preventing disease before it strikes. Eschewing the one-size-fits-all approach in favour of assessing individual circumstances, precision health encourages people to actively monitor and work towards improving their health so as to lower the risk of disease [2]. Personal health monitoring is part of this vision, allowing people to not only increase understanding of their health but also to receive recommendations for any necessary interventions. Significant advances in the Internet of Things (IoT) over the last decade has led to the rapid rise of wearable sensors, which are increasingly being used for health monitoring outside traditional clinical settings. Wearable sensors can collect and measure physiological data such as vital signs, which can be combined with health records and questionnaires to determine lifestyle habits and medical history. Additionally, ambient sensors can monitor environmental factors such as air quality and weather, which have a significant impact on health.
There are several crucial issues affecting sensor-based personal health monitoring systems, which can be distilled into seven key challenges. The first of these is interoperability. Heterogeneous sensor observations, differing data transmission technologies, and disparate standards for describing health data all contribute to interoperability issues in personal health monitoring systems. Additionally, the representation of health domain knowledge and its integration with sensor data remains a challenging task [3]. By its nature, sensor data is dynamic and complex, necessitating interpretation into higher-level concepts or situations [4]. Situation analysis involves the use of sensor data to detect the current state of a given environment (situation detection), while anticipating possible future states (situation prediction) [5]. The representation of domain knowledge is essential for facilitating situation analysis from sensor data and supporting subsequent decision-making. The decision support process augments human judgement, assisting clinicians in navigating complex medical decisions [6] and supporting patients in making informed health decisions outside clinical settings [7]. Both the situation analysis and decision support processes must incorporate context awareness. Dey and Abowd [8] define context as any information that can be used to characterize the situation of an entity, including location, identity, activity, and time. Such information is essential for accurate situation analysis and targeted decision support. Moreover, since the health domain is a high-stakes one, explainability is gaining traction as a pivotal aspect of AI-driven health systems [9]. Finally, given the probabilistic nature of health outcomes and the limitations of sensor and other health data, there is inherent uncertainty in the situation analysis and decision-making processes [10]. Thus, effective uncertainty handling is critical in sensor-based personal health monitoring systems.
Semantic Web technologies, which are widely used in the health domain, can alleviate some of these key challenges. The goal of this study is to systematically map the state of the art in the use of Semantic Web technologies in sensor-based personal health monitoring systems. The contributions of this paper are as follows:
  1. 1.
    We present a systematic mapping of the field based on 43 systems that are systematically selected as representative of the current state of the art.
  2. 2.
    We critically evaluate the extent to which the systems address the seven key challenges, i.e. interoperability, context awareness, situation detection, situation prediction, decision support, explainability, and uncertainty handling. We discuss the role and limitations of Semantic Web technologies in managing each challenge.
  3. 3.
    We undertake a quality assessment of the selected systems based on the data and devices used, system and components development, rigour of evaluation, and accessibility of research outputs.
  4. 4.
    Following an analysis of the current architectures, components, functionalities, and development tools, we propose a reference architecture to provide guidance for the design and development of new systems.
  5. 5.
    We highlight inadequacies in existing systems and outstanding issues in the field, thereby identifying potential directions for future research.
The remainder of this paper is structured as follows. Section 2 provides an overview of personal health monitoring using sensors and highlights how Semantic Web technologies can enhance sensor-based health monitoring systems. Section 3 discusses related reviews and surveys, motivating the novelty and importance of this study. Section 4 details the methodology used to conduct the study, including the search strategy and the inclusion and exclusion criteria, culminating in a summary of the selected systems. Section 5 discusses the seven key challenges that such systems must address, and critically analyses the capacity of the systems to deal with these challenges, while Section 6 analyses the quality of each system. The architectures of the selected systems are discussed in Section 7 and a reference architecture is proposed. Section 8 summarises the main findings of the study, discusses its limitations, and makes recommendations for future research directions. Finally, Section 9 concludes the study.
2 Background
2.1 Sensor-based personal health monitoring
Sensors used for health monitoring are typically worn, implanted, or placed in close proximity to the human body. When several such sensors are used at the same time, they form a wireless body sensor network (BSN), also known as a body area network (BAN) [11]. This is part of the IoT paradigm, in which sensor-based “things” connect and exchange data over a shared network such as the Internet. Two categories of physiological data can be collected from health monitoring sensors: vital signs and biological signals (biosignals). The primary vital signs are heart rate, blood pressure, respiratory rate, temperature, and blood oxygen saturation [12]. Biosignals are space- or time-based records produced from electrical, chemical, or mechanical activity within the body during a biological event such as a beating heart [13]. They include records of electrical activity in the body such as electrocardiograms (ECG) for the heart, electromyograms (EMG) for the skeletal muscles, and electroencephalograms (EEG) for the brain, as well as data from photoplethysmography (PPG), an optical sensing technology consisting of an LED and a photodetector to detect blood volume changes [14]. In addition to physiological data, physical activity data such as daily step count can also be captured by sensors. This data provides important contextual information about an individual’s lifestyle, which can enhance health monitoring.
Health monitoring sensors are generally either wearable or implantable. Wearable sensors are worn on the body or are otherwise integrated with clothes and shoes. Such sensors include electrodes for measuring electrical signals, thermal sensors for measuring temperature, and PPG sensors. Smart watches and bands are the most commonly used wearable sensors, but earables (devices placed in the ear) have recently emerged as a promising alternative [15]. In contrast, implantable sensors operate from within the human body. Although they are much less commonly used than wearable sensors, they are particularly useful for monitoring chronic illness as well as post-surgery monitoring to minimise complications and avoid readmission [16]. Health monitoring sensors also include portable devices that can measure physiological and activity data but cannot be practically worn or used for prolonged periods of time. Examples of these include blood pressure monitors and pulse oximeters, as well as smartphones which contain sensors such accelerometers, which measure acceleration, and gyroscopes, which measure orientation and angular velocity [17]. Additionally, ambient sensors are increasingly being incorporated in health monitoring to monitor the state of the external environment, such as temperature, humidity, and air quality, as these factors have a significant impact on human health [12, 18].
Sensors, while essential for personal health monitoring, also contribute significantly to the identified challenges. The heterogeneity of sensor devices, observation data, and measurement procedures can hinder interoperability in personal health monitoring systems [19]. The dynamicity and complexity of sensor data requires expert knowledge to interpret and analyse it. This affects both situation analysis and decision support. Furthermore, sensors can contribute to uncertainty. Data is uncertain when the degree of confidence about what is stated by the data is less than 100% [20]. This can arise when there is missing data or when all the relevant attributes cannot be measured by the available sensors [21]. Some of these challenges can be addressed by the incorporation of Semantic Web technologies.
2.2 Semantic Web technologies
Three overlapping Semantic Web technologies have emerged as the most prominent over the years: ontologies, knowledge graphs, and linked data [22].
2.2.1 Ontologies
Arguably, the key technology underpinning the Semantic Web is ontologies, which have been widely used for reasoning and representation in sensor-based systems [4]. Their ability to represent knowledge formally and unambiguously not only enhances interoperability but is also useful in capturing the domain knowledge necessary for situation analysis and subsequent decision support. Several ontologies have been developed to support the description of sensors and their observations, which is critical in any sensor-based system. Two particularly prominent sensor ontologies are the semantic sensor network (SSN) ontology [19] and the Smart Appliances REFerence (SAREF) [^1] https://saref.etsi.org ontology [23]. Both are standardised ontologies developed by the World Wide Web Consortium (W3C) and the European Telecommunication Standardization Institute (ETSI) respectively, with the aim of enabling semantic interoperability. However, while SSN was developed for sensors and sensor-based systems in general, SAREF focuses on smart appliances and IoT devices. The latest version of SSN is based on the Sensor, Observation, Sample, and Actuator (SOSA) ontology [24], which provides it with a lightweight, user-friendly, and extendable core. SAREF has mappings to SSN, from which it borrows modelling patterns for several classes [23, 25].
As domain-agnostic ontologies, both SSN and SAREF require augmentation to meet application-specific requirements [26]. SAREF provides a suite of ontologies that extend the core ontology for different domains, including two that are relevant for personal health monitoring: SAREF4EHAW [^2] https://saref.etsi.org/saref4ehaw/ for eHealth and ageing well and SAREF4WEAR [^3] https://saref.etsi.org/saref4wear/ for wearable devices. SAREF4EHAW provides support for modelling concepts such as health system actors (including patients and caregivers) and health devices (including wearables), with the wearable concept linked to the SAREF4WEAR ontology. An additional extension, SAREF4Health, was developed to address the limitations of SSN and SAREF in representing real-time ECG time series data exchanged between mobile devices and cloud gateways [25]. In contrast to SSN, SAREF is targeted at industry developers rather than ontology experts [25], making it practical for real-world applications. Furthermore, its extensions for the health domain provide a solid foundation for building semantic personal health monitoring systems. Additional representational support can be obtained by integrating resources such as standardized clinical terminologies and medical knowledge bases.
2.2.2 Knowledge graphs
A knowledge graph can generally be understood as a knowledge base of real-world data represented in a graph-based data model. Ontologies are a vital building block of knowledge graphs, used to define the data schema (such as properties, restrictions, and relationships) and enable semantic reasoning and entailment [27]. Knowledge graphs have seen increasingly widespread use in the health domain. Their graph structure enables the conceptualisation, representation, and integration of data [27]. This is advantageous in health monitoring systems, where the integration of various sources of health data is critical. An example of this is the Precision Medicine Knowledge Graph (PrimeKG), which integrates diverse biomedical data from multiple sources with the goal of enabling precision medicine analyses [28]. Previous research has also explored the automatic construction of knowledge graphs from electronic health records [29, 30], which can then be used for clinical decision support. Additionally, knowledge graphs have been proposed for health risk prediction [31], drug discovery [32], and as a tool for explainability in AI-driven health systems [33, 34]. Knowledge graphs have also proven useful in sensor-based systems, for example by providing graph-based visualisations of the data generated by IoT devices, which can then be queried in real time [35].
2.2.3 Linked data
Both knowledge graphs and ontologies can be published using a linked data approach [22], whereby uniform resource identifiers (URIs) are used to identify distinct resources [36]. When the emphasis is on free use, modification, and sharing, it is referred to as Linked Open Data [22]. Linked data has been proposed for augmenting and representing sensor data in order to improve its accessibility and interoperability [37]. In the health domain, it has been explored in applications ranging from drug discovery [38] to the representation of electronic health records [39]. Linked data can contribute to interoperability by ensuring heterogeneous health data is stored in a consistent format and structure. However, its use in health monitoring is not well explored in the literature.
2.2.4 Languages and standards
The development of Semantic Web technologies is facilitated using different languages and standards. Resource Description Framework (RDF) [^4] https://www.w3.org/RDF, a standard for the description and exchange of interconnected data in the form of subject-predicate-object triples, can be considered one of the core building blocks of the Semantic Web. Several extensions to RDF have been proposed. These include RDF Schema (RDFS) [^5] https://www.w3.org/wiki/RDFS, which provides a vocabulary to enrich RDF data; RDF-star [^6] https://w3c.github.io/rdf-star/cg-spec, which allows the subject or object of a triple to refer to another triple; and the Notation 3777https://w3c.github.io/N3/spec specification, which extends the representational abilities of RDF by supportive declarative programming and allowing the access of online knowledge. Other important standards in the Semantic Web community are: eXtensible Markup Language (XML), a markup language and file format; Web Ontology Language (OWL) [^8] https://www.w3.org/OWL, a language for constructing ontologies; Semantic Web Rule Language (SWRL) [^9] https://www.w3.org/Submission/SWRL, a language for expressing rules and logic; Shapes Constraint Language (SHACL) [^10] https://www.w3.org/TR/shacl, a language for describing RDF graphs, which also includes a rules language; and SPARQL Protocol and RDF Query Language (SPARQL) [^11] https://www.w3.org/TR/rdf-sparql-query, a language for retrieving and manipulating RDF data. SPARQL-star extends SPARQL to allow querying and updating of RDF-star data.
3 Related reviews
Several reviews related to sensors, Semantic Web technologies, and the health domain have been published. These reviews can generally be categorised into three overlapping groups, which are illustrated as a Venn diagram in Figure 1. The reviews in Group 1 focus on the use of Semantic Web technologies in the health domain; those in Group 2 review the use of sensors and IoT in the health domain; those in Group 3 review the use of Semantic Web technologies with sensor and IoT data; and finally, Group 4 consists of other related reviews that do not fit neatly into any of the first three groups. The related reviews are discussed in detail in the remainder of this section and summarised in Table 1.
Figure 1: Venn diagram illustrating the three focus areas of this study as well as the different groups of related reviews.
3.1 Group 1: Semantic Web technologies in the health domain
This group of reviews explores the use of Semantic Web technologies in healthcare. Zenuni et al. [40] review ontologies and semantic data repositories used in different aspects of the health domain, including hospital systems and health datasets. A similar review is conducted by Haque et al. [9], who explore themes such as e-healthcare, disease diagnosis, and information management. Peng et al. [41] and Hammad et al. [42] focus on semantic approaches for health data integration and management, including data from wearable devices. Dimitrieski et al. [43] review ontologies and ontology alignment approaches in healthcare, while Jing et al. [44] focus on ontologies for rule management in clinical decision support systems. Although the reviews in this group provide a good overview of the ways in which Semantic Web technologies have been used in the health domain, half of them do not mention sensors or IoT at all, while the other half do not include sensor data as a major focus.
3.2 Group 2: Sensors and IoT in the health domain
This group considers the use of sensors and IoT in the health domain. Islam et al. [45] and Yin et al. [46] conduct general surveys on IoT for healthcare, covering a broad range of considerations on the topic including networks, communication standards and protocols, and cybersecurity. The review by Qi et al. [47] focuses on the use of IoT in personalised healthcare systems, including sensor devices and data processing techniques. Philip et al. [48] explore advances in the field such as cloud computing, while Albahri et al. [49] focus on health monitoring systems for telemedicine applications, highlighting techniques that support the connection of hospital services to remote patients. There have also been reviews specifically focusing on the state of the art in wearable sensors for health monitoring, such as those by Cusack et al. [18], Dias and Cunha [12], and Majumder et al. [50]. Kim et al. [51] hone in on biosensors that detect biofluids, such as sweat and tears, while Baig et al. [52] highlight the potential of remote monitoring systems for clinical adoption. Punj and Kumar [53], Banaee et al. [54], and Andreu-Perez et al. [16] explore advances in wearable sensor data collection, mining, and processing, and Dang et al. [55] focus on statistical analysis and machine learning (ML) as modelling tools. While these reviews provide useful analyses on the role of sensors and IoT in health monitoring, they either do not mention Semantic Web technologies or do so briefly without an in-depth analysis of their role in health monitoring.
3.3 Group 3: Semantic Web technologies for sensors and IoT
This group reviews the intersection between Semantic Web technologies and sensors without being limited to a particular domain. Honti and Abonyi [56] and Rhayem et al. [57] explore the use of ontologies in IoT-based systems in different domains. Bajaj et al. [58] adopt a similar focus on ontologies, reviewing both general sensor ontologies as well as domain-specific ones for IoT. Compton et al. [59] present a review of the semantic specification of sensors using ontologies, analysing the range and expressive power of sensor ontologies. The review by Harlamova et al. [60] explores the challenges in the use of Semantic Web technologies in IoT, while Ye et al. [61] review the application of Semantic Web technologies in pervasive and sensor-driven systems. Although these reviews highlight the use of Semantic Web technologies with sensors and IoT, they are not specific to the health domain.
3.4 Group 4: Other reviews related to AI and technology in the health domain
A small number of reviews take a broader lens and consider different aspects of AI and technology in the health domain. This includes the concept of Healthcare 4.0, a term referring to the increasing digitisation of the healthcare industry. The reviews by Tortorella et al. [62] and Jayaraman et al. [63] broadly cover Healthcare 4.0, and highlight health monitoring systems that use IoT and sensors. However, only the review by Jayaram et al. [63] mentions ontologies and other knowledge representation techniques. A recent review by Kumar et al. [64] on AI in healthcare mentions IoT and knowledge graphs, but neither of these are the focus of the review. The review by Behera et al. [65] focuses on techniques used to create healthcare systems modeled on human cognitive processes such as perception and thought. They highlight cognitive IoT as a future research direction through wearable sensors, while also mentioning Semantic Web technologies for knowledge representation. However, neither the Semantic Web technologies nor sensors are discussed in detail.
3.5 Summary
Table 1 summarises the related reviews. The current study differs from existing work by focusing on the use of sensors and Semantic Web technologies for personal health monitoring, with both sensor data and Semantic Web technologies being primary points of focus. Additionally, the majority of the related reviews and surveys do not take a systems perspective, whereas this study highlights how the different system components are integrated and discusses the development methodologies and tools, evaluation approaches, and architectures of the included systems.
Table 1: Summary of related reviews and their focus areas.
Group Review Year Semantic Web technologies Healthcare/ health monitoring Sensors/IoT
1. Semantic Web technologies in the health domain Dimitrieski et al. [43] 2016
Hammad et al. [42] 2020
Haque et al. [9] 2022
Jing et al. [44] 2023
Peng et al. [41] 2020
Zenuni et al. [40] 2015
2. Sensors and IoT in the health domain Albahri et al. [49] 2018
Andreu-Perez et al. [16] 2015
Baig et al. [52] 2017
Banaee et al. [54] 2013
Cusack et al. [18] 2024
Dang et al. [55] 2023
Dias and Cunha [12] 2018
Islam et al. [45] 2015
Kim et al. [51] 2019
Majumder et al. [50] 2017
Philip et al. [48] 2021
Punj and Kumar [53] 2019
Qi et al. [47] 2017
Yin et al. [46] 2016
3. Semantic Web technologies for sensors and IoT Bajaj et al. [58] 2017
Compton et al. [59] 2009
Harlamova et al. [60] 2017
Honti and Abonyi [56] 2019
Rhayem et al. [57] 2020
Ye et al. [61] 2015
4. Other related reviews Behera et al. [65] 2019
Jayaraman et al. [63] 2020
Kumar et al. [64] 2023
Tortorella et al. [62] 2020
This study 2024
- the area is a main focus area of the review - the area is partially addressed, but is neither discussed in depth nor a focus area of the review - the area is not addressed at all in the review
4 Methodology
4.1 Objectives and reporting strategy
In order to achieve our goal of mapping the state of the art in the use of Semantic Web technologies in sensor-based personal health monitoring systems, the following are the objectives of this study:
  1. 1.
    To systematically select systems that represent the state of the art in the use of Semantic Web technologies in sensor-based personal health monitoring systems.
  2. 2.
    To determine the extent to which the seven key challenges are addressed by the selected systems.
  3. 3.
    To determine the role that Semantic Web technologies play in addressing these challenges.
  4. 4.
    To highlight inadequacies in existing systems and provide recommendations for future research.
Given the goal and objectives of this work, a mapping study was the most appropriate approach. Although systematic mapping studies are similar to systematic literature reviews in terms of the systematic process of searching for and selecting studies, literature reviews aim at synthesizing evidence while mapping studies structure a research area through classification and categorisation in order to discover research trends [66]. The study was conducted and is reported using the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) [67] framework. To further ensure the quality of the study, we adhered to the following quality assessment criteria as described by Kitchenham et al. [68]:
  1. 1.
    “The inclusion criteria are explicitly defined in the paper”: The inclusion and exclusion criteria are specified in Section 4.3.
  2. 2.
    “The authors have either searched four or more digital libraries and included additional search strategies or identified and referenced all journals addressing the topic of interest”: Six digital libraries were searched, and additional records were identified by using the preliminary search results and related reviews to search for similar studies. More details on the search strategy are given in Section 4.2.
  3. 3.
    “The authors have explicitly defined quality criteria and extracted them from each primary study”: The systems are analysed in Section 5 based on the seven identified challenges and the criteria is outlined in Table 8. Additionally, the quality of each system is assessed and discussed in Section 6 based on criteria outlined in Table 13.
  4. 4.
    “Information is presented about each paper so that the data summaries can clearly be traced to relevant papers”: A summary of all the included systems is shown in Table 4, with all systems fully cited. A GitHub repository [^12] https://github.com/mbithenzomo/semantic_phms_mapping_study has been created for this study, which includes copies of the relevant papers and other supplementary material.
4.2 Search strategy
Six digital libraries were searched between 9th and 12th February 2024: ACM Digital Library [^13] https://dl.acm.org, IEEE Xplore [^14] https://ieeexplore.ieee.org/Xplore, PubMed [^15] https://pubmed.ncbi.nlm.nih.gov, ScienceDirect [^16] https://www.sciencedirect.com, Scopus [^17] https://www.scopus.com, and Web of Science [^18] https://www.webofscience.com. For all libraries, abstracts, titles, and/or keywords were searched using terms related to the topic of the study, at the intersection of five areas: Semantic Web technologies, sensors, the health domain, monitoring, and systems. The search strings used are shown in Table 2. Boolean operators were used for a more specific search, although the ScienceDirect library had a limit on the number of Boolean operators that could be used per search. This library also did not allow the use of wildcard characters. Across all libraries, the results were filtered to only include literature published in or after 2012 to ensure a state of the art study. Additionally, where possible, the results were filtered to only include conference papers and journal articles published in English. This filtered out other types of literature such as surveys and reviews, books and book chapters, research abstracts, posters and conference proceedings, as well as articles written in languages other than English.
Table 2: Search strings used in digital library search.
Area Search strings
Semantic Web technologies semantic*, ontolog*, knowledge graph, linked data
Sensors sensor*, iot, internet of things, wearable*, device*, body area network
Health domain health*, medic*
Monitoring monitor*, track*, remote, tele*, distributed, continuous, daily
Systems system, framework, application, architecture
Rayyan [69], an online tool for the management of systematic reviews, was used to facilitate the screening process. The initial digital library search yielded 960 records, which were then screened and assessed for eligibility. The resulting records, together with the related review articles discussed in Section 3, were then used to identify further potentially relevant studies. This was done using two online tools, Semantic Scholar [^19] https://www.semanticscholar.org and Connected Papers [^20] https://www.connectedpapers.com. The supplementary search results were filtered by publication year (in or after 2012), and the titles and abstracts were screened to identify records containing the relevant search terms. 62 additional records were identified through the supplementary search process, resulting in a total of 1,022 identified records.
4.3 Inclusion and exclusion criteria
This study includes only peer-reviewed journal articles and conferences papers in English. Only systems that incorporate semantic techniques were included for analysis, and those without a well-defined semantic technique as an integral component were excluded. Additionally, because a system consists of several integrated components, studies reporting the development of only one component (for example, an ontology) were excluded. Of particular interest are sensors that measure physiological data (that is, biosignals and vital signs) and/or physical activity data (for example, daily step count). Applications of sensors outside health monitoring , such as activity recognition, fitness, or nutrition, were excluded. Furthermore, systems that do not have an analysis, inferencing, or reasoning component were also excluded. These inclusion and exclusion criteria are summarised in Table 3.
Table 3: Inclusion and exclusion criteria.
# Criteria Inclusion Criteria Exclusion Criteria
C1 Publication year The year of publication is 2012 or later. The year of publication is earlier than 2012.
C2 Language The publication is written in English. The publication is written in a language other than English.
C3 Publication type The publication is a peer-reviewed journal article or conference paper reporting original research. The publication is either not peer-reviewed (e.g. research abstracts, posters, books, and keynotes), is a collection of works (e.g. conference proceedings), or does not report original research (e.g. reviews, surveys, and position papers).
C4 Accessibility The publication is open access or can otherwise be accessed by the authors, e.g. through institutional access. The publication cannot be accessed without additional payment.
Table continued on next page.
Inclusion and exclusion criteria.
# Criteria Inclusion Criteria Exclusion Criteria
C5 Multiple integrated components The publication must report on a system, framework, application, or architecture consisting of several integrated components. Studies reporting the development of only one component (e.g. an ontology).
C6 Semantic Web technologies The system incorporates Semantic Web technologies as an integral component. Semantic Web technologies are either poorly defined or do not form an integral component of the system.
C7 Health monitoring The system focuses on health monitoring. The system has a focus outside the health domain, or is related to health but does not focus on health monitoring (e.g. systems focusing solely on other areas such as activity recognition, sports, fitness, and nutrition).
C8 Sensors for physiological and/or physical activity data The system incorporates sensors that measure physiological data (i.e. biosignals and vital signs) and/or physical activity data (e.g. daily steps). The system does not incorporate sensors or the sensors incorporated do not measure physiological or physical activity data.
C9 Analysis & Reasoning The system has an analysis, inferencing, or reasoning component. The system does not analyse or reason over the sensor data.
C10 Extended work If the system has been extended in later work, the more recent version is included in the review. The system is extended in later work.
4.4 Selection results
From the 1,022 identified records, 303 duplicates were removed resulting in 719 unique records. Next, preliminary screening was done by reviewing the title and abstract of each record. At this stage, records were excluded for reasons such as not being focused on the health domain or not involving health monitoring. We also found that a number of records had bypassed some of the filters that were applied in the initial identification stage, such as publication year and language. 533 records were excluded based on the title and abstract screening. The remaining 186 papers were read in full to determine if they still met the inclusion criteria. One reason for exclusion at this stage was if the system had been extended in later work and the extension was one of the systems being assessed. In such cases, the extension was included in the study while the previous work was excluded. Additionally, a small number of publications were excluded due to the full text being inaccessible without additional payment. Ultimately, 43 systems were selected for inclusion in this study. Figure 2 shows a PRISMA flow diagram illustrating the identification, screening, eligibility, and inclusion process.
Figure 2: PRISMA flow diagram outlining the selection process.
4.5 Summary of selected systems
A summary of the 43 selected systems is shown in Table 4, while Figure 3 shows the distribution of the systems according to the publication year. The year of publication ranges from 2012 to 2024, with 2021 being the most common. In terms of the application area, 25 focus on a particular disease or diseases, while the remaining 18 provide a solution for general health monitoring. Regarding the types of Semantic Web technologies used in the systems, nearly all of them make use of ontologies. The exceptions are the systems proposed by Yu et al. [70] and Zhou et al. [71] which use only knowledge graphs, and the one proposed by Xu et al. [72], which uses both linked data and a knowledge graph. Similarly, Reda et al. [73] use both linked data and an ontology, while Stavropoulos et al. [74] and Zafeiropoulos et al. [75] use both a knowledge graph and an ontology. Table 4 also provides an overview of other complimentary technologies and techniques used, as well as the architecture type of each system. These aspects are discussed in more detail in Sections 5 and  7 respectively.
With respect to the types of data used in the systems, all 43 systems collect physiological and other body data. 13 systems additionally incorporate environmental data from ambient sensors, while 19 consider data from existing health and medical records. Table 5 provides an overview of the types of sensor and non-sensor data used in the systems. A detailed analysis of the data sources, including sensor devices and existing datasets, is provided in Section 6.1.
Figure 3: Bar graph showing the distribution of the systems by year of publication.
Table 4: Summary of systems selected for this study.
# System Year Application Semantic Technologies Other Technologies & Techniques Architecture Type
1 Akhtar et al. [76] 2022 Parkinson’s Ontology Agents; CDL; Rules Layered; multi-agent
2 Ali et al. [77] 2021 Diabetes; ABP Ontology ML; NLP; Rules Layered
3 Ali et al. [78] 2020 Heart disease Ontology ML; Rules Layered
4 Ali et al. [79] 2018 Diabetes Ontology FL; Queries; Rules Layered
5 Alti et al. [80] 2022 Diabetes Ontology Agents; Queries; Rules Layered; multi-agent; service-oriented
6 Chatterjee et al. [81] 2021 Obesity Ontology Queries; Rules Modular
7 Chiang and Liang [82] 2015 General monitoring Ontology FL; Rules Modular
8 De Brouwer et al. [83] 2022 Headache disorders Ontology ML; Queries Modular
9 El-Sappagh et al. [84] 2019 Diabetes Ontology Queries; Rules Modular
10 Elhadj et al. [85] 2021 General monitoring Ontology Rules Layered
11 Esposito et al. [86] 2018 Arrhythmia Ontology FL; Rules Layered
12 Fenza et al. [87] 2012 General monitoring Ontology Agents; FL; Rules Layered; multi-agent; service-oriented
13 Garcia-Valverde et al. [88] 2014 General monitoring Ontology ML; Rules Unclear
14 Hadjadj and Halimi [89] 2021 General monitoring Ontology Queries; Rules Layered
15 Henaien et al. [90] 2020 General monitoring Ontology ML; Queries; Rules Layered
16 Hooda and Rani [91] 2020 Diabetes; heart disease Ontology Queries; Rules Modular
17 Hristoskova et al. [92] 2014 Heart failure Ontology Rules Service-oriented
18 Hussain and Park [93] 2021 Stroke Ontology ML; Queries; Rules Modular
[…]
21 Khozouie et al. [96] 2018 General monitoring Ontology Rules Modular
22 Kim et al. [97] 2014 General monitoring Ontology Queries; Rules Layered
23 Kordestani et al. [98] 2021 Kidney disease; skin disease Ontology ASP; Rules; BN Layered
24 Lopes de Souza et al. [99] 2023 Hypertension Ontology Queries; Rules Layered; modular
25 Mavropoulos et al. [100] 2021 General monitoring Ontology Agents; ML; NLP; Rules Layered; modular; single-agent
26 Mcheick et al. [101] 2016 Stroke Ontology BN Layered
27 Mezghani et al. [102] 2015 Diabetes Ontology ML; Queries; Rules Layered; service-oriented
28 Minutolo et al. [103] 2016 Arrhythmia Ontology FL; Rules Modular
29 Peral et al. [104] 2018 Diabetes Ontology ML; NLP; Rules Unclear
30 Reda et al. [73] 2022 General monitoring Linked Data; Ontology Queries; Rules Layered
31 Rhayem et al. [105] 2021 Gestational diabetes Ontology Queries; Rules Modular
32 Spoladore et al. [106] 2021 Diabetes; pulmonary disease Ontology Queries; Rules Layered
33 Stavropoulos et al. [74] 2021 Multiple sclerosis Knowledge graph; Ontology Rules Modular
34 Titi et al. [107] 2019 General monitoring Ontology Queries; Rules Layered
35 Vadillo et al. [108] 2013 General monitoring Ontology Agents Layered; multi-agent
36 Villarreal et al. [109] 2014 Diabetes Ontology None specified Layered
37 Xu et al. [72] 2017 General monitoring Linked data; Knowledge graph CBR; Queries Layered; service-oriented
38 Yu et al. [70] 2022 Paediatric asthma Knowledge graph ML; NLP; Rules Modular
39 Yu et al. [110] 2017 General monitoring Ontology Queries; Rules Layered
40 Zafeiropoulos et al. [75] 2024 Parkinson’s Knowledge graph; Ontology ML; Queries; Rules Modular
41 Zeshan et al. [111] 2023 General monitoring Ontology Queries; Rules Modular
42 Zhang et al. [112] 2014 General monitoring Ontology Rules Layered; modular
43 Zhou et al. [71] 2022 General monitoring Knowledge graph ML Modular
ABP - Abnormal blood pressure; ASP - Answer set programming; BN - Bayesian network; CBR - Case-based reasoning; CDL - Contextual defeasible logic; FL - Fuzzy logic; ML - machine learning; NLP - Natural language processing.
Table 5: Sensor data and other types of data used in the systems.
# System Supported sensor data and other types of data
1 Akhtar et al. [76] Body (BP, HR, BT, ECG, EEG, EMG); Ambient (temperature, CO & CO2 levels, motion)
2 Ali et al. [77] Body (BP, SpO2, BT, HR, ECG, EEG, BG); Other (health/medical records, social networks, smartphone)
3 Ali et al. [78] Body (RR, SpO2, BP, BT, HR, EMG, EEG, ECG, BG, cholesterol, position, activity); Other (health/medical records)
4 Ali et al. [79] Body (ECG, EEG, EMG, HR, BP, BG, cholesterol, range of motion)
5 Alti et al. [80] Body (HR, BG, motion); Other (GPS)
6 Chatterjee et al. [81] Body (BP, BG, activity); Ambient (temperature, humidity); Other (interviews, questionnaires, weather forecast, health/medical records)
7 Chiang and Liang [82] Body (BP, HR, BG, cholesterol); Ambient (motion, indoor & outdoor temperature, humidity)
8 De Brouwer et al. [83] Body (Acceleration, HR, blood volume pulse, galvanic skin response, skin temperature); Other (daily headache diary, questionnaires)
9 El-Sappagh et al. [84] Body (BP, HR, BG); Other (health/medical records)
10 Elhadj et al. [85] Body (BT, HR, BP, RR, SpO2); Ambient (temperature, humidity, location, motion); Other (health/medical records)
11 Esposito et al. [86] Body (BT, HR, SpO2, acceleration)
12 Fenza et al. [87] Body (HR, BP, BT, SpO2, BG); Ambient (temperature)
13 Garcia-Valverde et al. [88] Body (HR, acceleration, orientation/angular velocity, magnetoresistance)
14 Hadjadj and Halimi [89] Body (BP, HR, BT, BG); Other (vehicle sensor data)
15 Henaien et al. [90] Body (SpO2, BP, HR, RR, BT); Ambient (temperature, light, motion); Other (health/medical records)
16 Hooda and Rani [91] Body (BP, HR, BG, ECG); Other (health/medical records)
17 Hristoskova et al. [92] Body (BP, HR, SpO2, ECG); Other (health/medical records, WiFi location tag)
18 Hussain and Park [93] Body (ECG); Other (health/medical records)
[…]
21 Khozouie et al. [96] Body (BP, BT, SpO2, ECG, EMG, acceleration, orientation/angular velocity); Ambient (temperature, humidity, CO & O2 levels); Other (GPS)
22 Kim et al. [97] Body (BP, other unnamed vital signs); Ambient (temperature, illumination, humidity, wind); Other (weather forecast, news, weather indices)
23 Kordestani et al. [98] Body (BT, other unnamed vital signs); Ambient (temperature); Other (health/medical records)
24 Lopes de Souza et al. [99] Body (HR, BP, BT, acceleration, orientation/angular velocity); ; Other (user-submitted data)
25 Mavropoulos et al. [100] Body (BP, BG, sleep); Ambient (video); Other (health/medical records)
26 Mcheick et al. [101] Body (BP, blood flow velocity)
27 Mezghani et al. [102] Body (BP, HR, BG); Other (health/medical records)
28 Minutolo et al. [103] Body (BT, HR, SpO2, acceleration)
29 Peral et al. [104] Body (BG); Other (the web, existing databases, health/medical records)
30 Reda et al. [73] Body (HR, BT, BP, weight, calories burned, step count); Other (self-reported data)
31 Rhayem et al. [105] Body (BT, BP, HR, BG, cholesterol, activity); Ambient (temperature, humidity); Other (health/medical records)
32 Spoladore et al. [106] Body (HR, SpO2)
33 Stavropoulos et al. [74] Body (HR, step count, sleep)
34 Titi et al. [107] Body (BT, BP, HR, BG); Ambient (temperature, humidity)
35 Vadillo et al. [108] Body (HR, BT, BP, SpO2, BG); Ambient (motion, temperature, occupancy of bed / chair, CO levels)
36 Villarreal et al. [109] Body (BP, BT, BG)
37 Xu et al. [72] Body (BP, ECG, BG); Other (health/medical records)
38 Yu et al. [70] Body (BP, HR, sleep, exercise, weight); Other (health/medical records, self-reported data)
39 Yu et al. [110] Body (HR, BP, body fat); Other (mobile applications)
40 Zafeiropoulos et al. [75] Body (HR, movement data, sleep); Other (health/medical records)
41 Zeshan et al. [111] Body (BT, BP, HR); Other (GPS)
42 Zhang et al. [112] Body (BP, BT, HR, SpO2); Other (health/medical records)
43 Zhou et al. [71] Body (BP, HR, RR, BT, SpO2, BG, uric acid, cholestrol, lipoproteins, triglycerides, sleep); Ambient (inhalable particulate matter, CO2, temperature, formaldehyde, total volatile organic compounds); Other (health/medical records)
BG - blood glucose; BP - blood pressure; BT - body temperature; CO2 - carbon dioxide; CO - carbon monoxide; ECG - electrocardiogram; EEG - electroencephalogram; EMG - electromyogram; GPS - global positioning system; HR - heart rate; O2 - oxygen; RR - respiratory rate; SpO2 - blood oxygen saturation
5 Challenges in health monitoring systems
This section examines the role of Semantic Web technologies in addressing the seven key challenges identified in Section 1, as well as the contribution of other complimentary technologies and techniques that are incorporated into the systems. Additionally, a critical evaluation is provided assessing the extent to which each system succeeds in addressing these key challenges. Although there is a broader range of challenges facing sensor-based health monitoring systems, we have necessarily had to delimit the scope of this article. By focusing our analysis on these seven salient challenges, we aim to provide an in-depth assessment of how effectively they have been addressed in the current state of the field. However, we briefly discuss some other considerations, including privacy and security as well as usability, at the end of the section.
5.1 Interoperability
Interoperability can be defined as the ability of different components or systems not only to exchange information but also to make use of it [113]. There are three types of interoperability identified in the health domain: technical, semantic, and process interoperability [114, 113]. Technical interoperability refers to the way data or information moves from one system or component to another. Related to this is syntactic interoperability, which provides a structure and syntax for the transmitted data [115]. Semantic interoperability refers to the ability of the recipient to understand and make use of the received data, whereas process interoperability concerns the way in which different systems are used in actual work settings. A subset of this is clinical interoperability, through which patients can be seamlessly transferred between different care teams [113].
5.1.1 Technical interoperability
Differing data transmission technologies can contribute to a lack of technical interoperability in health monitoring systems, particularly those that use a range of different sensors. Data transmission protocols used in sensors include Bluetooth, Bluetooth Low Energy, ANT+, and Zigbee, with the first three being the most common among wearable devices today [11]. Interoperability among these different protocols can be achieved using gateway devices, which receive data from different sensors and transmit it to cloud services [116]. This is done by Ali et al. [79], who use a router as a gateway to receive sensor data and transmit it to the internet. A number of the systems [78, 80, 84, 85, 93, 96, 99, 104, 109, 112] use a mobile phone as a gateway device or base station, typically receiving sensor data via Bluetooth or Bluetooth Low Energy and transmitting it to the cloud via Wi-Fi or mobile data.
5.1.2 Syntactic interoperability
While technical interoperability is associated with hardware components and infrastructure, syntactic interoperability is usually associated with data formats [117]. There are several standards that are widely used to promote syntactic interoperability among systems. Among them is the ISO/IEEE 11073 standard, which provides a common format for communication involving medical devices and patient health data, with an emphasis on vital signs. This is used by El-Sappagh et al. [84] for message formatting between body sensors and the base unit. Other important standards for health data are provided by Health Level 7 (HL7). One of these is Fast Health Interoperability Resources (FHIR), which describes data formats, resources, and an application programming interface (API) through which health information can be exchanged [113]. El-Sappagh et al. [84] integrate FHIR in their proposed system, converting sensor data from the ISO/IEEE 11073 standard to FHIR resource formats. Additionally, the system receives data in FHIR format from hospital information systems. In this way, both sensor data and data from hospital systems are in the same format. FHIR resources can be defined using different data formats [^21] https://build.fhir.org/resource-formats.html, including XML, JavaScript Object Notation (JSON), and Terse RDF Triple Language (Turtle).
5.1.3 Semantic interoperability
The next type of interoperability is semantic interoperability, which is concerned with the meaning of the exchanged information. Semantic interoperability can be achieved through the use of unambiguous codes and identifiers, which can be provided by existing standard classifications and terminologies [113]. Ontologies are, of course, a well-established way to embed semantic interoperability in a system [118]. Within the medical domain, many existing medical terminologies are available as ontologies, including SNOMED CT [^22] https://bioportal.bioontology.org/ontologies/SNOMEDCT, the International Classification of Diseases (ICD) [^23] https://bioportal.bioontology.org/ontologies/ICDO, and the International Classification for Nursing Practice (ICNP) [^24] https://bioportal.bioontology.org/ontologies/ICNP. Among the systems, SNOMED CT is the most commonly used [81, 84, 98, 99, 73, 105, 107, 71]. ICNP is used by Elhadj et al. [85] and Henaien et al. [90] , while ICD is used by Spoladore et al. [106] and Yu [70] (ICD-11, the latest version) as well as Titi et al. [107] (ICD-10). The Unified Medical Language System (UMLS) [119] is a large thesaurus that integrates multiple terminologies of medical knowledge. It is used by Peral et al. [104] and Zhou et al. [71]. Another thesaurus is Medical Subject Headings (MeSH), which is used for indexing, cataloguing, and searching health information, and is integrated in the system proposed by Reda et al. [73]. Spoladore et al. [106] incorporate the International Classification of Functioning, Disability and Health (ICF) [^25] https://icd.who.int/dev11/l-icf/en.
[…]
Semantic Web technologies also provide a means to represent sensors and the data they capture. Sensors can be represented with varying degrees of expressiveness. Concepts that can be captured about sensors include unique identifier, manufacturer, location of deployment, dimensions, operating conditions, type of data captured, and hierarchy with regard to related sensors [59]. Similarly, various sensor data concepts can be represented, such as the property being observed, units of measurement, and measurement timestamps. Most systems represent sensor and sensor data concepts in ontologies, with 13 reusing and extending existing sensor or device ontologies, namely SSN/SOSA [81, 84, 85, 94, 105, 74, 107], SAREF and its extensions [83, 89, 99, 75], the Amigo device ontology [92], and the Moving Objects ontology [105]. The reuse of existing sensor ontologies, particularly established ones such as SAREF, can contribute to a higher degree of expressiveness for sensor and sensor data concepts. This is because these validated ontologies provide rich modelling of such concepts, facilitating more effective querying of and reasoning on sensor data, which is essential for situation analysis. Comprehensive sensor ontologies also support sensor management, allowing sensors to be catalogued based on their attributes as captured in ontologies [59].
Foundational ontologies can contribute to semantic interoperability by providing unambiguous and domain-independent concept definitions [120]. Three of the selected systems directly incorporate a foundational ontology. El-Sappagh et al. [84] use the Basic Formal Ontology, while De Brouwer et al. [83] and Stravropoulos et al. [74] use the DOLCE+DnS (Description and Situation) Ultra Lite (DUL) ontology. Other systems indirectly integrate foundational ontologies by reusing other ontologies that have already incorporated them. For example, the SSN ontology uses DUL as its upper ontology [19], and the SAREF ontology also has an indirect reference to DUL through its mappings to the SSN ontology [121]. Consequently, any system that reuses the SSN or SAREF ontologies inherits an indirect connection to DUL.
5.1.4 Process interoperability
The final type of interoperability is process interoperability, which focuses on how systems and components work seamlessly together in real-world settings. One way to enhance process interoperability in health monitoring systems is through the integration of sensor data with comprehensive health and medical records [114]. The inclusion of these records allows health monitoring systems to complement and extend healthcare provided in clinical settings. Health and medical records provide additional information that is useful for health monitoring, such as an individual’s disease history, laboratory test results, medications taken, allergies, and previous hospital admissions. 19 systems integrate existing records in some way, with most of them represented using ontologies. The systems proposed by Ali et al. [78, 79], El-Sappagh [84], and Rhayem et al. [105] have the most comprehensive records, capturing laboratory tests, prior disease diagnoses, and lifestyle information such as exercise, nutrition, alcohol consumption, and smoking status. Some systems use medical records to extract diagnosis status [77], while others use them to extract an individual’s risk factors for disease [78]. These records can also be used to overcome limitations of sensor data such as missing values, as was done by Ali et al. [77]. Besides health and medical records, data from social networks and other web and mobile applications can also be used to complement sensor data. For instance, Ali et al. [77] use social networking data for monitor individuals’ mental health through sentiment analysis.
5.1.5 The role of Semantic Web technologies
Semantic Web technologies are critical in the achievement of semantic interoperability among the selected systems. Many of the systems make use of terminologies such as SNOMED CT and ICD through ontologies, and access knowledge graphs such as DrugBank and Diseasome which are published as linked data. This allows the systems to access and reason with standardised health domain knowledge. Additionally, many systems use ontologies and knowledge graphs to represent sensor data and health records, which can enhance process interoperability. Semantic Web technologies can also contribute to syntactic interoperability, albeit in an indirect capacity. For instance, El-Sappagh et al. [84] map FHIR resources to an ontology, allowing for interoperability between their proposed system and hospital information systems that use FHIR. However, Semantic Web technologies do not inherently provide support for technical interoperability since they operate at a higher, more abstract level to formally represent and derive meaning from the data. Therefore, in order to achieve technical interoperability, health monitoring systems must leverage data transmission protocols and devices.
5.2 Context awareness
An important aspect of health monitoring is the ability to take context into consideration, which is critical for situation analysis since contextual information enhances sensor data and supports its interpretation. Consider a case where an individual’s heart rate is suddenly elevated. If the individual is engaged in exercise, the increased heart rate is expected. However, if the individual is at rest, this could be a cause for alarm. Therefore, health monitoring systems must be able to adapt based on the context of the individual being monitored. The four most common aspects of context are location, time, identity (of a person or agent), and activity (or events) [122, 123].
5.2.1 Location
Ye et al. [122] highlight three types of locations that can be represented: symbolic locations, coordinate locations, and regions. The systems proposed by Akhtar et al. [76], Chiang and Liang [82], and Vadillo et al. [108] keep track of the different rooms in a house where an individual may be, while those proposed by Khozouie et al. [96], Titi et al. [107], and Zeshan et al. [111] indicate more generally the place the monitored individual is (for example, “home” or “hospital”). These are symbolic locations. One purpose of such locations is to allow the systems to suggest relevant services based on the type of space currently occupied, as is the case in the system proposed by Chiang and Liang [82]. In the system proposed by Hristoskova et al. [92], the clinician’s location (i.e. the room they occupy in a hospital) is used to determine which device to send notifications to, optimising for the closest device. Likewise, Zeshan et al. [111] determine the closeness between the monitored individual and their caregivers in order to select which caregiver or clinician to notify in case of abnormal sensor observations. This is similar to the system proposed by Alti et al. [80], which supports a GPS sensor that captures the current coordinates of the monitored individual. In this system, location is used to select devices closest to the user from which to deploy health services so as to increase efficiency and minimise inter-device communication costs. Coordinate locations also serve the purpose of alerting caregivers and emergency services of the exact location of a person in the event of a medical emergency, as is suggested by Rhayem et al. [105]. The system proposed by Hadjadj and Halimi [89] integrates health monitoring in the public transport system, and therefore includes location sensors in public transportation vehicles. The final type of location is regions, which are geometrical two- or three-dimensional representations of locations [122]. This type of location is used in the system proposed by Kim et al. [97] in order to advise users of region-specific situations, such as adverse or dangerous weather. Similarly, El-Sappagh et al. [84] use the spatial region class from the Basic Formal Ontology to represent the patient’s current location, as well as the placement of the sensors. Despite the importance of location as an aspect of context, less than half of the systems include it.
5.2.2 Time
In contrast, nearly all of the systems include the concept of time. Observation time is the most common way time is incorporated in the systems, with many systems capturing the exact timestamp for each sensor observation [80, 81, 83, 84, 85, 86, 89, 93, 95, 99, 100, 103, 105, 74, 107, 108, 75, 111]. Besides observation time, the time at which certain events occur can be captured, for example calls to emergency services [80]. This allows the systems to display or analyse trends over time. Additionally, Alti et al. [80] capture the time intervals in which reports should be sent. Rather than a timestamp, several systems also capture the general time of day during which observations or activities occur. For instance, Ali et al. [77] divide the time at which daily activities are done into morning, afternoon, and evening. Peral et al. [104] use mealtimes as a point of reference, which is particularly important when taking blood glucose measurements. They distinguish between pre-breakfast, pre-lunch and pre-dinner readings.
Duration and frequency are other important aspects of time. Duration can be captured for physical activity [81, 106], sleep [81, 74, 71], symptoms [101], and treatment [107, 72], and disease [75]. De Brouwer et al. [83] capture headache duration as well as the duration of events that influence headaches, such as stress and sleep. Symptom duration can influence the risk for certain illnesses, while specifying treatment duration ensures medication reminders are sent only during the prescribed period. When combined with thresholds, duration can be useful in identifying different situations. For example, Stavropoulos et al. [74] determine that an individual has a lack of movement if they have fewer than 500 steps and their heart rate has been less than 100 beats per minute for longer than 800 minutes. Frequency is used by Chiang and Liang [82], Spoladore et al. [106], and Yu et al. [110] as a metric for physical activity. Mezghani et al. [102] capture the frequency of sensor observations, while Villarreal et al. [109] capture the frequency of detected diseases.
[…]
5.2.3 Identity
Identity, which pertains to the actors in a system, is another important aspect of context [122]. This includes the definition of individuals and their properties, such as name, address, gender, and age. For health monitoring, this can include additional information such as weight, height, and blood group. This is the most ubiquitous aspect of context in the systems, with nearly every system including personal information about the monitored individuals. Besides personal properties, identity also encompasses different user roles within the system. Most of the systems [76, 77, 78, 79, 80, 81, 83, 84, 85, 86, 89, 91, 92, 93, 94, 95, 98, 99, 100, 101, 104, 105, 106, 74, 107, 109, 70, 75, 111] support different users besides the individual being monitored, typically including clinicians and in some cases, caregivers and family members. Identity also includes agents, which are used in the agent-based systems [76, 80, 94, 95, 100, 108]. Agents [^27] An agent is a computer system situated in some environment that is capable of acting autonomously in order to achieve some goal(s) [125] have been applied extensively in the health domain [126] as well as in sensor-based systems [127]. The agent-based approach offers several advantages. For example, agents can be used as personal assistants to support humans in performing tasks and services [128]. This is explored in the system proposed by Mavropoulos et al. [100], which includes a smart virtual agent that clinicians can interact with via voice commands. Similarly, Yu et al. [70] implement an AI chatbot to answer user questions. Agent-based architectures and their advantages are discussed in greater detail in Section 7.
5.2.4 Activity
[…]
5.2.5 Other types of context
Besides location, time, identity, and activity, other types of contextual information are incorporated in the systems. Alti et al. [80] include hardware and network information as part of context, such as available communication protocols, CPU speed, battery power, and memory size. This information is used to ensure the efficient deployment of health services. Similarly, Zeshan et al. [111] use battery level and device response time to determine which caregiver or clinician’s device to send notifications to. Hristoskova et al. [92] incorporate media devices and their properties in their interpretation of context. For example, the screen size of devices such as mobile phones and tablets is used to determine how to display the health monitoring results. For small screens, the results are summarised. An important factor in health monitoring is the state of a person’s environment. A number of the systems use environmental data such as temperature and humidity from ambient sensors to provide additional context [76, 82, 85, 87, 96, 97, 98, 105, 107, 108, 71]. Weather data sources such as forecasts and indices are used by Kim et al. [97] to supplement sensor data, while the systems proposed by Akhtar et al. [76], Khozouie et al. [96], Vadillo et al. [108], and Zhou et al. [71] include sensors to monitor air quality by checking the levels of gases such as carbon monoxide, carbon dioxide, and oxygen. Contextual information can also include details about an individual’s diet, medication, and emotional state. These details are collected in the system proposed by De Brouwer et al. [83] through self-reporting via a mobile app. Similarly, Zafeiropoulos et al. [75] capture information about the user’s fatigue level.
5.2.6 The role of Semantic Web technologies
Ontologies appear to be particularly useful for the representation of contextual information, and majority of the selected systems reuse existing ontologies to do so. For instance, OWL-Time [^28] https://www.w3.org/TR/owl-time, an ontology that describes temporal properties of real-world objects such as sensors, is reused by a number of systems [83, 100, 105, 107, 110]. Friend of a Friend (FOAF) [^29] http://xmlns.com/foaf/0.1, an ontology that describes people profiles, is also widely reused among the systems [85, 90, 100, 73, 106, 107, 110]. Additionally, sensor ontologies, while not focused solely on contextual information, also include some aspects of context. For example, SAREF and SSN/SOSA ontologies include timestamps for sensor observations. Knowledge graphs can also be used to represent contextual information, as is done by Yu et al. [70], who use a knowledge graph to capture patient profile data from health records. Table 6 summarises the contextual information included in the systems and indicates which types of contextual information are captured using Semantic Web technologies.
Table 6: Summary of contextual information captured in the systems and represented using Semantic Web technologies.
# System Types of contextual information Represented using Semantic Web technologies
1 Akhtar et al. [76] L (patient); T (temporal logic); I (profile, user roles); O (air quality, weather) L; I; O
2 Ali et al. [77] T (activity); I (profile, user roles); A (step count, intensity level) None
3 Ali et al. [78] I (profile, user roles); A (intensity level) I; A
4 Ali et al. [79] I (profile, user roles); A (range of motion, intensity level) I; A
5 Alti et al. [80] L (patient, device); T (observation timestamps; report intervals); I (profile, user roles); O (hardware info, network info) L; T; I; O
6 Chatterjee et al. [81] T (observation timestamps, activity duration, sleep duration); I (profile, user roles); A (step count, intensity level, exercise type); O (weather) T; I; A; O
7 Chiang and Liang [82] L (patient); T (activity, exercise frequency); I (profile); A (detection, motion); O (weather, illumination) L; T; I; A; O
8 De Brouwer et al. [83] L (patient, headache location); T (observation timestamps, trigger duration); I (profile, user roles); A (sleep, physical activity, commute); O (diet, medication, mood) L; T; I; A; O
9 El-Sappagh et al. [84] L (patient, sensor); T (observation timestamps); I (profile, user roles); A (intensity level) L; T; I; A
10 Elhadj et al. [85] L (patient); T (observation timestamps); I (profile, user roles); O (weather) L; T; A; O
11 Esposito et al. [86] T (observation timestamps); I (profile, user roles); A (step count, intensity level) T; I; A
12 Fenza et al. [87] I (profile); O (weather) None
13 Garcia-Valverde et al. [88] T (situation timestamps); I (profile); A (recognition, intensity level) T; I; A
14 Hadjadj and Halimi [89] L (vehicles, bus stop); T (observation timestamps); I (profile, user roles); O (passenger count, vehicle status) L; T; I; O
15 Henaien et al. [90] L (patient); I (profile); A (motion); O (weather) L; I; A
16 Hooda and Rani [91] I (profile, user roles) I
17 Hristoskova et al. [92] L (clinician, device); T (risk horizon); I (profile, user roles); O (device size) L; I; A; O
18 Hussain and Park [93] T (observation timestamps; time-domain features); I (profile, user roles) None
[…]
21 Khozouie et al. [96] L (patient); T (observation timestamps & intervals); I (profile); A (type); O (air quality, weather) L; T; I; A; O
22 Kim et al. [97] L (patient’s region); I (profile); O (weather) L; A; O
23 Kordestani et al. [98] T (episode timestamps); I (profile, user roles); O (weather) I; O
24 Lopes de Souza et al. [99] T (observation timestamps); I (profile, user roles); A (movement) T; I; A
25 Mavropoulos et al. [100] T (observation timestamps; time-domain features); I (profile, user roles); A (recognition) T; I; A
26 Mcheick et al. [101] T (symptom duration); I (profile, user roles) I
27 Mezghani et al. [102] T (observation start/end date, observation frequency, anomaly timestamps); I (profile, user roles) T; I
28 Minutolo et al. [103] T (observation timestamps); I (profile); A (step count) T; I; A
29 Peral et al. [104] T (observation timestamps); I (profile, user roles) T; I
30 Reda et al. [73] L (patient); T (observation timeframe); I (profile, user roles); A (step count, type, intensity) L; T; I; A
31 Rhayem et al. [105] L (patient, device trajectory); T (observation timestamps); I (profile, user roles); A (recognition); O (weather) L; T; I; A; O
32 Spoladore et al. [106] T (exercise timestamps, duration & frequency); I (profile, user roles); A (exercise type) T; I; A
33 Stavropoulos et al. [74] T (observation timestamps, sleep duration, time taken to fall asleep); I (profile, user roles); A (sleep quality, step count, intensity level) T; A
34 Titi et al. [107] L (patient); T (observation timestamps, intervals, & duration); I (profile, user roles); A (type, intensity level); O (weather) L; T; I; A; O
L - location; T - time; I - identity; A - activity; O - other Table continued on next page.
Summary of contextual information captured in the systems and represented using Semantic Web technologies.
# System Types of contextual information Represented using Semantic Web technologies
35 Vadillo et al. [108] L (patient); T (observation timestamps); I (profile); A (detection); O (air quality, weather) L; I
36 Villarreal et al. [109] T (disease duration & frequency); I (profile, user roles); A (type) T; I; A
37 Xu et al. [72] T (treatment duration); I (profile, user roles) I
38 Yu et al. [70] I (profile, user roles); A (exercise); O (diet, medication) I; A; O
39 Yu et al. [110] L (patient); T (disease progression, medical event timestamp; exercise frequency); I (profile); A (type) L; T; I; A
40 Zafeiropoulos et al. [75] T (disease duration, observation timestamp); I (profile, user roles); A (gait analysis, sleep quality); O (fatigue) T; I; A; O
41 Zeshan et al. [111] L (patient, caregiver, clinician); ; T (observation timestamps); I (profile, user roles); O (battery level; response time) L; T; I
42 Zhang et al. [112] T (observation timestamps); I (profile) I
43 Zhou et al. [71] T (movement timestamps, sleep duration); I (profile, user roles); A (sleep quality); O (air quality, weather) None
L - location; T - time; I - identity; A - activity; O - other
5.3 Situation analysis
A situation can be understood as a higher-level interpretation of sensor data that is relevant and of interest in an application domain [4]. Personal health monitoring systems should be capable of situation analysis. This entails both the detection and the prediction of health situations, which constitute two of the seven key challenges. We discuss each of them in turn.
5.3.1 Situation detection
In health monitoring systems, situation detection can take a variety of forms. One of these is the categorisation of individual sensor observations based on whether they are within or outside a given range as determined by domain knowledge. For example, in the system proposed by Akhtar et al. [76], when vital signs such as temperature and heart rate are outside the normal range, the situation is classified as an emergency. Likewise, Elhadj et al. [85] classify expected observations as normal, while observations outside the normal ranges are classified as abnormal. They also include a third classification, wrong, for faulty observations from malfunctioning sensors. Similar threshold-based situation categories are used in many of the systems [80, 83, 88, 89, 92, 94, 96, 99, 104, 105, 107, 109, 75, 111, 112]. Thresholds have also been used to classify physical activity based on level of intensity [81, 84, 86, 88, 94]. A better approach than using individual sensor observations is to consider different observations and personal attributes to classify individuals. This is done by Ali et al. [79], who classify the patient health condition as either healthy, moderate, or serious based on multiple sensor outputs and properties such as sex, weight, and height. Similarly, Chiang and Liang [82] classify situations as either healthy, moderate, or severe based on age, blood pressure, blood glucose, heart rate, and cholesterol.
[…]
With regard to techniques for situation detection, most of the systems implement some form of rule-based reasoning. Rules provide a way to implement expert knowledge in an if-then form, whereby if certain conditions are met, then a consequent conclusion is made or action taken. Despite their widespread use, rules have several limitations. Firstly, crisp rules are unable to handle uncertainty and ambiguity in sensor observations and the determination of health situations. To mitigate this, several systems incorporate fuzzy logic [79, 82, 86, 87, 103] and defeasible logic [76] in the rules. These techniques are discussed in greater detail in Section 5.6, which focuses on techniques for handling uncertainty in health monitoring. Secondly, rules are typically based on existing expert knowledge, and therefore cannot incorporate new knowledge that experts may be unaware of. Additionally, manually updating rules is time-consuming, making them difficult to scale. This challenge can be overcome using learned rules based on ML algorithms, which can acquire new, high quality knowledge automatically [130] and contribute to dynamic and adaptable rule-based systems. The systems proposed by Hussain et al. [93], Henaien et al. [90] and Peral et al. [104] extract rules from decision trees. However, caution should be exercised when using ML-derived rules, as they may still need verification and validation from domain experts. As an alternative to rule-based reasoning, Xu et al. [72] implement case-based reasoning, arguing that it is easier to capture human experiences using cases rather than rules. By searching for historical cases that are similar to the current case, their proposed system is able to obtain treatment plans that have been successful in the past.
In addition to the development of rules as discussed above, ML is also used in a number of the systems for the classification of diseases based not only on sensor data but also other data sources. Ali et al. [77] use a bidirectional long short-term memory (BiLSTM) model to detect diabetes and blood pressure, to classify sentiments from social networking data for mental health monitoring, and to classify drug side effects. Their proposed system uses domain ontologies to extract important features that can enhance the ML classification. Zhou et al. [71] also use a BiLSTM model for disease prediction. Other ML algorithms used include multi-layer perceptron for heart disease detection [78] and random forest for stroke detection [93]. ML is also used for physical activity classification, for example using the k-nearest neighbours [88, 100], decision trees [100], and random forest [94, 100] algorithms. Finally, ML can also be used to classify alert levels. This is done by Zafeiropoulos et al. [75], who use a graph neural network to distinguish between medium and high alerts. A full review of ML techniques for situation analysis in the health domain is outside the scope of this study. Readers are referred to the reviews by Ravì et al. [131] and Li et al. [132].
5.3.2 Situation prediction
A number of the selected systems explore the concept of risk as a situation prediction feature, since determining an individual’s risk profile for a certain condition can be used to predict future adverse health situations. For example, Alti et al. [80] use rules to determine the risk of death for diabetes patients based on high glucose levels and high heart rate. Rules are also used by Chiang and Liang [82] to determine the risk of arthritis recurring based on low temperatures and high humidity, and by Hristoskova et al. [92] also use rules to determine the risk of congestive heart failure (CHF) over a four-year time horizon. CHF risk stages are determined based on factors such as age, blood pressure, heart rate, and history of heart disease and diabetes. Similarly, in their use case of gestational diabetes, Rhayem et al. [105] use rules to determine the risk level for fetal loss based on age, presence of hyperglycemia, and presence of hypertension.
To support the identification of potential risks, future physiological readings can also be predicted using historic sensor observations, as is done by Peral et al. [104]. Their proposed system uses ML algorithms (support vector machine and logistic regression) to predict blood glucose levels over three-day and five-day windows. These predictions of sensor measurements can then be analysed to determine future health risks. Besides rules and ML, another technique used in some of the systems is Bayesian networks, which are probabilistic models in the form of directed acyclic graphs that can represent causal relationships among variables in a domain. Mchiek et al. [101] use a Bayesian network to calculate the risk of stroke occurring in the next seven days, based on risk factors such as age, presence of diabetes, high blood pressure, and symptom duration. This approach is also taken by Kordestani et al. [98] to determine the probability of the occurrence of kidney disease. The use of Bayesian networks and their contribution to explainability and uncertainty handling is discussed in greater detail in Sections 5.4 and  5.6 respectively.
As part of situation prediction, Zafeiropoulos et al. [75] explore the prediction of medication adherence using a graph neural network to identify potential lapses in pill dosing, while De Brouwer et al. [83] explore the detection of triggers as a means to anticipate potential headache attacks. Another useful aspect of situation prediction is the determination of the prognosis, i.e. expected progression, of a detected disease, although this is poorly explored in the systems. Hussain and Park [93] mention the intention to extend their system in future work to include automated stroke prognosis. In contrast, Yu et al. [110] include a disease progression class in their proposed ontology, representing past diagnoses or potential health risks and their associated times. However, the system does not include any methods to predict the progression of detected conditions.
5.3.3 The role of Semantic Web technologies
Semantic Web technologies can support situation analysis in two main ways: firstly, they formally represent important concepts and the relationships between them, i.e. sensor data, domain knowledge, contextual information, and even the situations themselves; and secondly, they support reasoning through which new knowledge can be derived from existing knowledge [4]. Although several situation-focused ontologies have been developed, including the Situation Theory Ontology [133] and the Scenes and Situations ontology [134], none of the selected systems extend any such ontologies. Despite this, Semantic Web technologies remain vital for situation analysis among the selected systems. Rule-based reasoning is the most common mechanism for situation analysis among the systems. These rules rely heavily on concepts that are formally defined in ontologies, and they are more often than not expressed in Semantic Web standard languages such as SWRL. The limitations of semantic-based approaches, such as scaling difficulties and inability to handle uncertainty, can be mitigated by combining them with complimentary techniques such as ML and Bayesian networks.
5.4 Decision Support
Decision support is the natural next step after situation analysis. Based on the detected and predicted situations, targeted support can be offered to mitigate adverse situations and promote favourable health outcomes.
5.4.1 Features and functionalities
Alerts and warnings are used in majority of the systems to warn of potentially dangerous situations and prompt mitigating action. These alerts are often sent to caregivers, clinicians, and emergency services depending on severity. However, several systems also take a patient-targeted approach, reminding the monitored individual about medications and exercise [82, 107, 108]. Alerts can also serve to remind users of medications, exercises, and medical tests. For example, the system proposed by Kordestani et al. [98] can remind clinicians to order additional laboratory tests when previously taken tests become out of date. A well-documented issue with alerts in the health domain is alert fatigue, a phenomenon in which users become desensitized to alerts due to their frequency [135]. Esposito et al. [86] mitigate this by differentiating between critical and non-critical abnormal situations, with the latter being sent out in a daily summary report email rather than an instantaneous notification for each case.
In addition to alerts, health monitoring systems can also trigger actions in response to adverse situations. For example, the system proposed by Alti et al. [80] triggers the injection of insulin in response to a blood glucose level above a certain threshold, while the system proposed by Hadjadj and Halimi [89] can trigger the opening of a vehicle door. Such systems must be integrated with an actuation device capable of carrying out the action. The system proposed by Titi et al. [107] includes several actuators such as a smoke alarm. Other systems are integrated with actuators capable of turning on lights and heaters [82], making emergency calls [105], or turning off water or gas if detected [108].
Another decision support feature is the generation of suggestions or recommendations for the mitigation of adverse situations. Several systems offer recommendations for lifestyle modifications, such as diet and exercise [78, 79, 81, 84, 97, 105, 106, 109, 70], as well as medication [84, 85, 92, 98, 104, 105, 107, 72]. An important factor when choosing appropriate treatment is the side effects of medications and how different medications interact with each other. Ali et al. [77] use drug review websites to collect data on side effects, while Elhadj et al. [85] keep track of medication interactions as well as patient allergies. This information assists clinicians in prescribing appropriate medications for each patient. Related to recommendations is the ability for the monitored individual to seek out relevant and trusted medical information. For example, the system proposed by Rhayem et al. [105] includes a notification module that allows patients to contact a clinician and receive recommendations and treatments from them.
5.4.2 Quality of decision support
An important factor in the quality of decision support is user agency. Rather than simply providing recommendations, decision support tools should allow decision-makers the agency to engage in decision-making by helping them to: 1.) identify and narrow down options; 2.) identify possible outcomes for each option; 3.) judge which outcomes are most likely; 4.) identify the value of each option based on their impact on stakeholders; 5.) make trade-offs using the aforementioned criteria; and finally, 6.) understand how and why the tools work [136]. This approach to decision support aligns with the human-centred AI paradigm, which advocates for AI systems to augment and enhance human capabilities and performance, rather than automating them away [137]. Considering the first five of these criteria, none of the selected systems demonstrate this level of decision support. Among the systems that provide intervention recommendations, none offer more than a single option for a particular situation, nor do they identify the possible outcomes of the recommendation. Thus, users are likely to either dismiss the recommended interventions or accept them blindly [136], neither of which is optimal. We discuss the sixth criterion in Section 5.5 on explainability.
Furthermore, the soundness of recommended interventions can be ensured by incorporating established and clinically validated medical guidelines. This can contribute to the acceptance of health monitoring systems by medical professionals and regulatory bodies. Despite this, few of the selected systems mention the use of such guidelines. They include Lopes de Souza et al. [99], who incorporate risk level classifications from the American Heart Association; De Brouwer et al. [83], who use the International Classification of Headache Disorders’ criteria to issue relevant alerts; and Ali et al. [78] and Hristoskova et al. [92], who the Framingham Risk Score to determine coronary heart disease risk.
5.4.3 The role of Semantic Web technologies
Similarly to situation analysis, the main value that Semantic Web technologies add to decision support in the selected systems is their ability to represent important concepts, particularly situations and domain knowledge, which can then be reasoned over. This is consistent with previous research findings on the use of Semantic Web technologies for decision support [138], with rules being the most common reasoning tool among the selected systems.
5.5 Explainability
A critical consideration in health monitoring systems is explainability, which contributes to the overall trustworthiness and adoption of such systems. For the purposes of this study, we adopt the perspective that explainability is essentially equivalent to interpretability [139, 140], which in turn can be defined as the degree to which a system’s operations can be understood by a human [141]. Two commonly used and complementary approaches to explainability in AI systems are: 1. prioritising human understanding of generated outputs; and 2. providing explicit explanations for those outputs [139].
5.5.1 Inherently interpretable techniques
One way to prioritise human understanding of system outputs is to use technologies and techniques that are inherently intuitive and easily comprehensible to humans. It can be argued that Semantic Web technologies meet this criteria and can contribute to the development of explainable systems; we discuss this in Section 5.5.3. However, the use of these technologies does not guarantee explainability. Other technologies and techniques implemented within the systems, and the ways in which they are combined, can also play a role in either enhancing or hindering the overall explainability of the system. For example, rules are inherently easy to understand [142], and nearly all the selected systems implement some form of rule-based reasoning. However, in some cases, rules are implemented for one aspect of the system while less interpretable techniques are used for other components. An instance of this is the system proposed by Ali et al. [78], where a deep learning model is used for disease prediction, while rule-based reasoning is applied for recommendation generation. This results in the decision support functionality being highly explainable, while the situation analysis component remains less so.
Bayesian networks can also be considered highly interpretable, as they can perform predictive and diagnostic reasoning [143] in a way that can be visually interpreted due to their graphical structure [144]. Predictive reasoning is done by Mcheick et al. [101], who use a Bayesian network to determine whether a person has a high risk of stroke based on risk factors. The reason for whether the risk is high or not can be traced back to the presence of risk factors. On the other hand, Kordestani et al. [98] use a Bayesian network for diagnostic reasoning, allowing for the understanding of a kidney disease diagnosis based on its causes. Fuzzy logic represents another class of interpretable techniques, since it allows variables and their classifications to be presented in a way that is intuitive [142]. Fuzzy approaches are used in a number of the selected systems; we discuss them in more detail in Section 5.6.1.
While ML is often criticized for its susceptibility to producing black box models, certain ML models are intrinsically interpretable, such as logistic or linear regression models and decision trees [145, 140]. However, even such models can be rendered uninterpretable as their complexity and scale increase, as is the case with large decision trees [145]. In cases where less interpretable ML models are used, they can be combined with Semantic Web technologies to enhance explainability. This is closely related to neuro-symbolic AI, in which the strengths of neural networks and symbolic methods are combined to achieve the best of both worlds [146]. A number of the selected systems explore such a hybrid approach, including Zafeiropoulos et al. [75] and Zhou et al. [71], who both use a knowledge graph to provide training data for deep learning models. Additionally, Ali et al. [77] use an ontology for feature extraction, providing some transparency into the selection of features for their BiLSTM model.
5.5.2 Post hoc explainability
The use of the use of inherently interpretable models should be prioritised in high-stakes domains [147]. Nonetheless, explaining the workings of a black box model after it has been trained is a well-established approach to explainability [140]. This can be done through methods such as the local interpretable model-agnostic explanations (LIME) technique [148], in which the black box model is approximated with an interpretable surrogate model that is then used to explain the original model’s predictions. However, such post hoc explainability methods are not mentioned in any of the selected systems.
Once the reasoning behind the system’s outputs has been determined, the next step is to communicate this to the users of the system. This explicit provision of explanations can also be considered a form of post hoc explainability. An explanation can be defined as the answer to a why-question [139], such as "Why was the situation classified as an emergency?", "Why was an alert sent to the clinician?", or "Why was this medication recommended?". A good explanation should have the ability to be contrastive (i.e. show the contrast between the situation chosen or decision made versus other possibilities), to distill a complex list of causes into one or two main ones, to highlight abnormalities, and to be general and probable [140]. The target audience is also an important consideration in the communication of explanations, since different audiences may value distinct aspects based on their role and perspective [149]. Health monitoring systems should therefore tailor explanations to the needs of various audiences, such as monitored individuals, clinicians, and caregivers. For certain audiences, like regulatory stakeholders, it may be more appropriate to include explanations as part of the overall system documentation. Only six of the selected systems report or indicate that some form of explanation is made available to users [76, 82, 85, 100, 57, 109, 70]. This is typically through a user interface, or, as in the system proposed by Akhtar et al. [76], as part of system logs. Chatbots and virtual agents, as implemented by Mavroppoulos et al. [100] and Yu et al. [70], can also provide explanations, since they are designed to answer queries from users.
5.5.3 The role of Semantic Web technologies
Explainability relies on domain knowledge, and Semantic Web technologies excel at structuring such knowledge formally and unambigiously [150]. Ontologies can contribute to the development of explainable systems from three perspectives: by providing sound and explicit knowledge reference models; by supporting common-sense reasoning through the representation of context-aware semantic information; and by facilitating flexible knowledge abstraction and refinement [150]. Since most of the systems use ontologies to represent expert knowledge and contextual information, this can be seen as a first step towards achieving explainability. Additionally, many ontology reasoners provide explanations of the reasoning process, although it is not clear whether these explanations are made available to the end users. Knowledge graphs can also contribute to better understanding of system outputs in several ways, including providing a graph-based visualisation of concepts, entity and relation extraction from unstructured data, enrichment of datasets, and inference and reasoning [33]. While the connection between linked data and explainability is not direct, the open accessibility and interconnection of knowledge shared using a linked data approach can nonetheless contribute to explainability.
5.6 Uncertainty handling
Given the uncertainty inherent in health decision-making as well as the high likelihood of ambiguity, noise, and missing values in sensor observations, health monitoring systems are greatly enhanced by being able to handle uncertainty. Despite this, only 19 of the systems addressed some aspect of this. The approaches used to handle uncertainty are summarised in Table 7 and discussed in detail in the remainder of this subsection.
Table 7: Approaches to handle uncertainty and the systems that use them.
Approach Systems
Fuzzy logic Ali et al. [79], Chiang and Liang [82], Esposito et al. [86]; Fenza et al. [87]; Minutolo et al. [103]
Bayesian networks Kordestani et al. [98], Mcheick et al. [101]
Answer set programming with probabilistic rules Kordestani et al. [98]
Defeasible logic Akhtar et al. [76]
Replacing or eliminating missing or invalid sensor data Ali et al. [77, 78]; Hooda and Rani [91]; Hussain and Park [93]; Reda at al. [73]; Rhayem et al. [105]; Titi et al. [107]
Filtering sensor data Ali et al. [77, 78]; Garcia-Valverde et al. [88]
5.6.1 Fuzzy logic
Fuzzy logic is a widely used technique for representing ambiguity and vagueness in sensor data [20]. It is used by five of the systems, making it the most commonly implemented uncertainty handling approach among the systems, besides the preprocessing of sensor data. In fuzzy logic, the truth of a statement is not binary (i.e. either true or false), but can rather be represented in a range from false to true. Therefore, rather than having crisp thresholds for different categories, fuzzy logic allows for values with different degrees of membership for the different categories. The process of converting crisp inputs into fuzzy sets is called fuzzification. For example, heart rate is represented in beats per minute, which can be classified into crisp categories. Generally, a heart rate greater than 100 beats per minute can be categorised as “fast”, a heart rate between 60 and 100 beats per minute can be categorised as “normal”, and a heart rate below 60 beats per minute can be categorised as “slow” [151]. However, with fuzzy logic, any given heart rate value has a certain degree of membership to any of the categories. For instance, a heart rate of 80 beats per minute may have a high degree of membership to the “normal” category (for example, 75%), a lower degree of membership to the “fast” category (for example, 20%), and an even lower degree of membership to the “slow” category (for example, 5%). Fuzzy logic provides a better approach to deal with boundary conditions. For example, when the heart rate is either 100 or 101, it can be reflected as mostly normal and to a lesser degree fast. Both Ali et al. [79] and Chiang and Liang [82] fuzzify sensor data such as blood pressure and heart rate, as well as attributes such as age and weight. Similarly, Esposito et al. [86] fuzzify the intensity of physical activity, which provides important context for heart rate thresholds. Fenza et al. [87] incorporate fuzzy logic with rules to determine the degree of membership to different situation categories based on different combinations of vital signs, while Minutolo et al. [103] use hybrid rules that incorporate both crisp and fuzzy variables. Fuzzy logic provides a simple but effective mechanism for representing imprecision and vagueness in sensor observations and allows this to be taken into account for more effective situation detection.
5.6.2 Bayesian networks
Bayesian networks are well known for managing uncertainty and have been widely used in the health domain [144]. Kordestani et al. [98] use a Bayesian network for probabilistic diagnosis of acute kidney injury. The Bayesian network models immediate (short-term) and background (long-term) causes of acute kidney injury, as well as its symptoms. They used experts to determine the conditional probabilities of the presence of acute kidney injury given these variables. Similarly, Mcheick et al. [101] represent risk factors for stroke using a Bayesian network.
5.6.3 Nonmonotonic reasoning
Monotonic reasoning holds that the rejection of an earlier conclusion must only be done if the evidence for the conclusion is also rejected. Contrastingly, nonmonotonic reasoning holds that an earlier conclusion can be rejected based on new evidence, even when earlier evidence was valid [152]. This ability to revise conclusions in the face of new evidence is useful in handling uncertainty. Defeasible logic is an example of nonmonotonic reasoning in which there are three kinds of rules: strict rules which can never have exceptions, defeasible rules which are typically true but can have exceptions, and undercutting defeaters which are weak possibilities [152]. Akhtar et al. [76] use defeasible logic to handle inconsistencies in sensor data as well as patient information. Another type of nonmonotonic reasoning is answer set programming (ASP), which is used by Kordestani et al. [98] to automatically customise treatments for each patient. They combine ASP with probability to reason with uncertain knowledge regarding treatment. Using probabilistic ASP rules, their proposed system obtains all possible treatment options for a medical episode and the associated probability of the episode occurring. If the probability of the episode occurring decreases with a particular treatment, then the treatment’s award value is increased. The treatment with the highest award value is ultimately selected by the system.
5.6.4 Preprocessing sensor data
Uncertainty can stem from various factors in sensor data, including ambiguous or imprecise readings, noise, or missing values caused by sensor malfunctions or network failures [21, 20]. Several systems have addressed the issue of missing values in sensor data. Ali et al. [77, 78] replace them with mean and median values from existing data, while Hooda and Rani [91] replace them using the preceeding value. Rhayem et al. [105] take the approach of removing any missing or unusual values, for example those outside the device measurement ranges. Similarly, Titi et al. [107] and Reda et al. [73] use rules to check whether sensor data falls within the expected minimum and maximum bounds. In their proposed system, Hussain and Park [93] use the Pan-Tompkins algorithm to detect the QRS complex in the ECG. This identifies beats without a QRS complex, which may be premature, missing, or ectopic, and are subsequently eliminated. To deal with noisy data, a few systems use filters to improve signal quality. Ali et al. [77, 78] use a Kalman filter to remove noise, while Garcia-Valverde et al. [88] use a moving average filter for the same purpose.
5.6.5 The role of Semantic Web technologies
Semantic Web technologies provide limited inherent support for uncertainty handling, but this can be mitigated through the techniques discussed in this section. Several extensions for Semantic Web standards have been proposed in the literature that make use of fuzzy logic and Bayesian inference, such as BayesOWL [153] and Bayes-SWRL [154], probabilistic extensions for OWL and SWRL respectively, as well as fuzzyDL [155], a fuzzy ontology reasoner. However, none of the selected systems that implement uncertainty handling techniques report using any such extensions, opting instead to define custom solutions. Another way that Semantic Web technologies can support uncertainty handling is through using ontology reasoners for inconsistency detection. Missing values or otherwise invalid data can be detected through reasoners such as HermiT [156] and Pellet [157] based on specified and inferred axioms; the use of these tools are discussed in greater detail in Section 6. Additionally, rule-based reasoning can be combined with ontologies to detect invalid data, an approach used by a few of the selected systems.
5.7 Other challenges
While we consider the seven challenges discussed above to be particularly salient in sensor-based personal health monitoring systems, we acknowledge that there are other factors that such systems must take into account. This subsection briefly discusses a few of them. As an in-depth analysis of these additional challenges is outside the scope of this study, we also include references to relevant articles that interested readers can consult.
5.7.1 Security and privacy
As health-related information is highly sensitive, security and privacy are important to consider. Particular aspects of this include security of data storage, network and transmission security, user authentication and access control, consent management, and the use of privacy-preserving techniques such as federated learning. For insights on security and privacy, we direct interested readers to the following articles: Rasool et al. [158] review security and privacy in the context of the Internet of Medical Things; Thapa and Camtepe [159] explore security and privacy challenges and techniques for health data in general; and finally, Kirrane et al. [160] provide an overview of security and privacy issues that relate to Semantic Web technologies.
5.7.2 Usability
Usability is another factor that health monitoring systems should consider, and can broadly be defined as the ease of use of a system [161]. It is a multi-faceted concept with several contributing factors, including understandability (which is closely related to explainability), attractiveness, and overall user satisfaction [161, 162]. Interested readers can refer to the following articles for more information: Saeed at el. [162] explore pertinent usability issues in health monitoring systems and identify possible solutions. With regard to the evaluation of usability, Maramba et al. [163] identify current methods used in usability testing in health monitoring applications, while Cho et al. [164] present a usability evaluation framework for mobile health applications. Finally, considering the usability of the sensors themselves, the reviews by Cusack et al. [18], Dias and Cunha [12], and Andreu-Perez et al. [16] highlight the types and characteristics of wearable sensors available for health monitoring.
5.7.3 Scalability
Scalability generally refers to the ability of a system to handle increased workload [165]. In the context of sensor-based health monitoring, this workload could arise from an increased number of sensors, other data sources, users, and services provided by the system. For further reading on scalability in IoT applications, readers can consult the review by Fortino et al. [166]. The reviews by Albahri et al. [49] and Phillip et al. [48] discuss scalability in the context of IoT and healthcare.
5.7.4 Ethics and regulatory compliance
The high-stakes nature of the health domain necessitates careful consideration of ethical issues. Although there are many benefits of technology-enabled personal health monitoring, there are also potential harms that it exposes. Many of these ethical issues overlap with the challenges already discussed, such as situation analysis (how accurate and reliable are the detected and predicted situations?), decision support (how appropriate are the suggested recommendations and how much autonomy do system users have?), explainability (to what extent can system outputs be understood?), and security and privacy (how secure is user data and how is consent managed?). An additional ethical concern is the cascade of care, a phenomenon in which incidental findings from screenings or monitoring result in further clinical care. Some of these ethical considerations can be enforced through regulation. Readers seeking further exploration on this challenge may consult the following articles: Morley et al. [167] comprehensively map the ethics of AI in healthcare; Kwan et al. [168] highlight the ethical issues associated with health monitoring applications; Nittari et al. [169] review ethical and legal challenges in telemedicine more broadly; Hassanaly and Dufour [170] explore the regulation of mobile health applications in the United States, the European Union, and France; and Thapa and Camtepe [159] explore legality and regulatory compliance from the perspective of security and privacy.
5.8 Summary
This section has provided an in-depth analysis of seven key challenges that must be addressed in health monitoring systems, as well as a brief discussion of additional challenges that are important in such systems. The role played by Semantic Web technologies in overcoming the seven key challenges has been critically examined. Additionally, non-semantic techniques that are incorporated in the systems have also been discussed. A full list of reused semantic resources, systems that reuse them, and the challenges they address can be found in Table 16 in the appendix. These resources include ontologies, knowledge graphs, and linked data, as well as vocabularies, taxonomies, and classifications. To summarise the section, we conduct an assessment of each system based on the seven key challenges. Finally, we discuss the results of the assessment, highlighting the challenges that are most neglected among the systems.
5.8.1 Challenges assessment
Table 8 summarises the different aspects related to the seven key challenges. While we consider these aspects to be highly important for achieving effective sensor-based personal health monitoring, it is possible that some of them may exceed the requirements for specific health conditions and may therefore not be essential in those particular cases. To assess the degree to which each system tackles the seven challenges, we use a rating scheme based on the identified aspects. A score of 1 point is assigned to each system for every aspect that is addressed. We use a five-point rating scale as follows:
  1. 1.
    ✗: None of the aspects are addressed by the system
  2. 2.
    Low: 40% or less of the aspects are addressed
  3. 3.
    Medium: between 41% and 69% of the aspects are addressed
  4. 4.
    High: between 70% and 90% of the aspects are addressed
  5. 5.
    Very High: more than 90% of the aspects are addressed
Table 8: Important aspects related to the seven key challenges.
Challenge Aspects
Interoperability 1. It is mentioned or illustrated how the system addresses the technical interoperability between the sensors and the rest of the system, e.g. using a gateway device, base unit/station, or established data transmission standards and protocols. 2. The system incorporates established standards or ontologies for describing sensor data, such as the SSN and SAREF ontologies. 3. The system makes use of established health and medical terminologies and nomenclatures such as SNOMED CT, ICD, and ICNP. 4. The system makes use of existing health data standards such as ISO/IEEE 11073, FHIR, and HL7 V2. 5. The system integrates existing health and medical records. 6. The system integrates other sources of data such as weather forecasts, social networks, and other web data.
Context awareness 1. The system includes and makes use of the concept of location, e.g. GPS coordinates, symbolic locations (“home”, “hospital”, “kitchen”), or geographic regions. 2. The system includes and makes use of the concept of time, e.g. observation timestamps, duration, etc. 3. The system includes different user roles, such as patient, caregiver, and clinician. 4. The system captures information related to an individual’s identity, such as name and address. 5. The system includes and makes use of the concept of activity, e.g. physical activity monitoring or activity recognition. 6. The system incorporates ambient sensor data in addition to physiological data from body sensors. 7. The system includes other types of contextual information, e.g. hardware and networking considerations.
Situation detection 1. The system can detect deviations or abnormalities in physiological measurements based on historical observations or known thresholds. 2. The system can classify individuals or situations into predefined categories, levels, or states related to health. 3. The system can detect medical conditions or diseases that are currently being experienced.
Situation prediction 1. The system can predict the risk of medical conditions, diseases, or other adverse effects in the future. 2. The system can predict future physiological measurements based on current or historical sensor observations. 3. The system can predict the prognosis of detected diseases.
Decision support 1. The system sends alerts and notifications for potentially dangerous situations. 2. The system sends reminders, e.g. for medication and exercise. 3. The system provides suggestions or recommendations for mitigation or treatment of adverse situations, e.g. medication, diet, or exercise. 4. The system provides decision support for more than one type of user, e.g. individuals, clinicians, caregivers, etc. 5. It is mentioned that the system incorporates established clinical practice workflows, medical guidelines, risk scores, scales, or information from medical or allied health professional bodies, the details of which are specified. 6. Rather than solely making recommendations, the system provides advanced support such as narrowing down options, identifying possible outcomes and their likelihood, and making trade-offs between options.
Explainability 1. The system uses inherently interpretable techniques as part of the situation analysis process (for example when classifying sensor data), or else applies a post hoc explainability method. 2. The system uses inherently interpretable techniques as part of the decision support process (for example when making recommendations), or else applies a post hoc explainability method.
Table continued on next page.
Important aspects related to the seven key challenges.
Challenge Aspects
3. It is mentioned or demonstrated that the system presents explicit explanations to the user for generated situations. 4. It is mentioned or demonstrated that the system presents explicit explanations to the user for generated decisions. 5. The presented explanations meet two or more of the criteria for good explanations as explained by Molnar [140], e.g. contrastiveness, conciseness, generalisability, and a focus on abnormalities. 6. The presented explanations are tailored to different target audiences (monitored individuals, clinicians, caregivers, regulators, etc), e.g. by providing different levels of detail or highlighting different aspects of the situation or decision.
Uncertainty handling 1. The system is able to handle uncertainty in the situation analysis process. 2. The system is able to handle uncertainty in the decision support process. 3. The system is able to handle missing, noisy, or otherwise invalid sensor data.
5.8.2 Discussion
Table 9 shows the number of systems with a particular rating for each challenge. Note that because situation detection, situation prediction, and uncertainty handling each have three aspects as per Table 8, a rating of “high” is not applicable based on the five-point rating scale. The combined radar chart in Figure 4 provides a visualisation of how well all the systems address the seven challenges. Separate radar charts for the individual systems are also available [^30] https://public.flourish.studio/visualisation/17843904, and the individual system ratings are shown in Table 17 in the appendix.
It is evident that more work is needed to address situation prediction, explainability, uncertainty handling, and to a lesser extent, interoperability and decision support. Situation prediction stands out as the most neglected challenge, with 30 of the 43 systems failing to address it altogether. This is closely followed by uncertainty handling, which 23 systems do not address. It is also notable that none of the selected systems have achieved a rating higher than medium for situation prediction, explainability, and uncertainty handling. Interoperability is not very well addressed among the systems, with most scoring a low rating for this challenge, although two of the systems achieve a high score. Moreover, although seven of the systems score either high or very high on decision support, most only moderately address it, and three do not address it at all. While nine systems score very highly on situation detection, the remainder score either a low or medium rating. However, all of the selected systems consider this challenge, making it generally well addressed. The same is true of context awareness, with 22 systems achieving a medium score, 17 achieving a high score, and one achieving a very high score.
Table 9: Counts of number of systems with each rating across the seven challenges.
Rating Challenge Interoperability Context awareness Situation detection Situation prediction Decision support Explainability Uncertainty handling
3 0 0 30 3 0 23
Low 25 3 12 12 19 35 15
Medium 13 22 22 1 14 8 5
High 2 17 N/A N/A 5 0 N/A
Very High 0 1 9 0 2 0 0
Figure 4: Combined radar chart showing the extent to which all the systems address the seven challenges.
6 System quality
In this section, we examine the quality of the selected systems as reported in the respective research articles. We consider four main criteria: the data sources and devices used to collect the data; the development methodologies and tools used; the evaluation approaches and rigour; and finally, the accessibility of research outputs. These factors all contribute to the credibility, reliability, and reproducibility of the reported systems. Additionally, this assessment can also be used to inform benchmarking for future research and development of such systems. We begin with a discussion of the methods of data collection, sources of data, and sensors reported in the systems. We then review the methodologies and tools used for the development of the different components of the system, including programming languages, libraries, frameworks, and other software. Next, we examine the evaluation approaches used to evaluate the system components and the systems as a whole. The last criteria we discuss is the accessibility of the resources and outputs of each system, including ontologies, data, code, and even user interfaces. We conclude by outlining the different aspects related to each criteria and then scoring the selected systems based on these aspects.
6.1 Data and devices
The data collection methodology varies among the systems. Many systems used existing datasets from publicly available repositories such as PhysioNet [^31] https://physionet.org and the University of California, Irvine (UCI) ML repository [^32] https://archive.ics.uci.edu. These systems and the datasets they use are summarised in Table 10.
Table 10: Existing health datasets used.
System Dataset Source
Ali et al. [77] Pima Indians diabetes dataset Multiparameter Intelligent Monitoring in Intensive Care (MIMIC-II) Drug review dataset UCI ML Repository PhysioNet UCI ML Repository
Ali et al. [78] Heart disease dataset (Cleveland, Hungary) UCI ML Repository
De Brouwer et al. [83] WESAD dataset Schmidt et al. [171]
Garcia-Valverde et al. [88] PAMAP2 Physical Activity Monitoring dataset UCI ML Repository
Hadjadj and Halimi [89] Vital signs of 15 Volunteers Figshare
Henaien et al. [90] Vital signs dataset University of Queensland
Hooda and Rani [91] Pima Indians diabetes dataset Heart disease dataset (Cleveland) UCI ML Repository UCI ML Repository
Table continued on next page.
Existing health datasets used.
[…]
Kordestani et al. [98] Chronic kidney disease dataset Dermatology dataset UCI ML Repository UCI ML Repository
Mavropoulos et al. [100] Heterogeneity Human Activity Recognition dataset CoNNL2003 dataset UCI ML Repository Sang et al. [172]
Peral et al. [104] Diabetes dataset Health Facts database UCI ML Repository Strack et al. [173]
Rhayem et al. [105] Various undisclosed datasets PhysioNet
Yu et al. [70] Undisclosed dataset Children’s Hospital, Zhejiang University School of Medicine
Zeshan et al. [111] Human vital signs dataset Kaggle
A smaller number of the systems used data collected from participants rather than existing data. For example, Ali et al. [79] collected data from 44 diabetes patients, while Esposito et al. [86] collected data from 10 healthy volunteers. Other systems that used this approach are those proposed by De Brouwer et al. [83], Hristoskova et al. [92], Hussain and Park [93], Stavropoulos et al. [74], Vadillo et al. [108], and Villarreal et al. [109]. Another approach was to simulate or manually generate the data. This was done by Alti et al. [80] who simulated temperature and camera data; Chatterjee et al. [81], who simulated the sensor, interview, and questionnaire data of four dummy participants; and Zafeiropoulos et al. [75], who simulated the sensor observations and health records of three virtual patients. Mchiek et al. [101] similarly generated 513 data records. Additionally, Stavropoulos et al. [74] simulated records in order to test the scalability of their proposed system. A significant number of the systems indicated the types of data and sensors supported by the systems, but did not mention the source of the data. It is unclear whether these systems were validated using actual sensor data, beyond a theoretical validation of the system functionality.
Less than half of the systems gave specific details of the devices used for data collection. While most of these were commercially available devices, the systems proposed by Kim et al. [97] and Lopes de Souza et al. [97] used custom-made prototypes. Table 11 indicates the types and descriptions of devices mentioned.
Table 11: Devices mentioned for data collection in the selected systems.
System Device type Name/description Data type
Chiang and Liang [82] Commercial Kinect Motion
De Brouwer et al. [83] Commercial Empatica E4 wristband Acceleration, HR, BVP, GSR, ST
Esposito et al. [86] Commercial Amiigo wristband Omron HJ-112 digital pocket pedometer BT, HR, SpO2 Acceleration
Hristoskova et al. [92] Commercial A&D UA-767PBT blood pressure monitor Nonin Avant 4000 digital pulse oximeter A&D UC-321PBT weight scale Welch Allyn Cardio Perfect 12 Lead ECG BP, HR SpO2 Body weight ECG
Hussain and Park [93] Commercial BioNomadix respiration (RSP) with ECG amplifier ECG patch from Life Science Technology Inc. ECG
[…]
Kim et al. [97] Prototype Smart wear consisting of upper body clothing made of flexible and stretchy material to which various sensors can be attached. To minimise the use of wires, a coin cell battery and wireless communication is used. BP, other unnamed vital signs
Table continued on next page.
Devices mentioned for data collection in the selected systems.
System Device type Name/description Data type
Lopes de Souza et al. [97] Prototype Various sensors (i.e. MPU-6050 gyroscope and accelerometer; MKB0805 HR and BP sensor; DS18B20 digital thermometer), are assembled on a T7 V1.3 MINI 32 ESP32 board which is then affixed using 3D printed plastic frame onto a bracelet. Acceleration, orientation, HR, BP, BT
Mavropoulos et al. [100] Commercial Samsung Galaxy S3 Mini smartphone LG Nexus 4 smartphone Video
Reda et al. [73] Commercial Fitbit Jawbone Step count, HR, calories burned
Spoladore et al. [106] Commercial Polar H1 chest strap COSMED E100 cycle ergometer ECG
Stavropoulos et al. [74] Commercial Fitbit Charge 3 Step count, sleep stages, HR
Vadillo et al. [108] Commercial Arduino e-Health Sensor Platform Tunstall Lifeline Connect+ home unit BG, HR, BT, BP, SpO2 Motion
Villarreal et al. [109] Commercial BodyTel Glucotel BG
Yu et al. [110] Commercial Fitbit Withings scale Step count Body weight
Zhang et al. [112] Commercial Equivital multi-parameter sensor BP, BT, HR, SpO2
BG - blood glucose; BP - blood pressure; BT - body temperature; BVP - blood volume pulse; ECG - electrocardiogram; GSR - galvanic skin response; HR - heart rate; SpO2 - blood oxygen saturation; ST - skin temperature
6.2 System and components development
6.2.1 Development methodologies
The use of a development methodology can streamline the process of developing Semantic Web technologies. In particular, the literature on ontology development methodologies is quite rich, with a large number of established methodologies proposed [174, 175]. There have also been several proposed approaches towards developing knowledge graphs [176] and ensuring the quality of linked data [177, 178]. Despite this, most of the systems did not report the use of a methodology in the development of the Semantic Web technologies. However, a small number of systems mentioned using a particular methodology. For example, Hadjadj and Halimi [89] used the NeOn framework [179], a scenario-based methodology for building ontologies, while Titi et al. [107] used an existing case-based ontology engineering methodology [180]. Although not a development methodology, Peral et al. [104] used the SemanTic Refinement of Ontology MAppings (STROMA) [181] approach for aligning corresponding concepts between different ontologies.
6.2.2 Development tools
Various languages, frameworks, and libraries were used to develop the systems. Among the systems that incorporated ontologies, Protégé [^33] https://protege.stanford.edu is most commonly cited as the ontology development platform of choice, mentioned in 22 of the systems. Protégé is an ontology editor that supports the latest OWL and RDF specifications. Another commonly used platform is Apache Jena [^34] https://jena.apache.org, a Java framework for building Semantic Web and Linked Data applications, mentioned in 16 of the systems. Both Protégé and Apache Jena are free and open source. When it comes to the Semantic Web languages, SWRL is the most commonly used rule language among the systems. However, Apache Jena includes a general purpose rule-based reasoner which is used by Chiang and Liang [82], Garcia-Valverde et al. [88], and Kim et al. [97]. Stavropoulos et al. [74] used SHACL to create rules, while Kordestani et al. [98] and Rhayem et al. [105] used Drools [^35] https://www.drools.org, a business rule management system. Programming languages can also be used to configure rules, as was done by Khozouie et al. [96] using Java. For queries, a majority of the systems used SPARQL, with some also using Apache Jena Fuseki, a SPARQL server, to publish their SPARQL endpoints. For storage, Mavropoulos [100] and Stavropoulos et al. [74] used GraphDB, while Spoladore et al. [106] used Stardog [^36] https://www.stardog.com. Both of these are enterprise semantic databases. Non-semantic database management systems were also used by a few of the systems, with MySQL being the most commonly cited.
6.3 Rigour of evaluation
[…]
A number of systems were evaluated based on quality of service metrics. For example, Esposito et al. [86] and Vadillo et al. [108] used the Architecture-Level Modifiability Analysis (ALMA) method to evaluate the potential costs associated with modifying their systems, such as by adding more sensors. Similarly, Alti et al. [80] evaluated their system based on execution time, optimality, application’s lifetime and number of discovered services. Yu et al. [70] evaluated their system using the Chronic Care Model (CCM), an established framework for chronic care management that includes criteria such as system design, self-management support, and decision support. The systems proposed by Hristoskova et al. [92], Kim et al. [97], Mavropoulos et al. [100], Stavropoulos et al. [74], and Villarreal et al. [109] were evaluated using user studies with patients or clinicians, with Likert scales typically used to scale the user feedback. Expert validation was also used to evaluate the systems, with the aim of ensuring maximum similarity between the system output and expert opinion. This approach was taken by Ali et al. [78], El-Sappagh et al. [84], Hadjadj and Halimi [89], Hristoskova et al. [92] and Khozouie et al. [96]. Additionally, a number of systems used query-based validation, where the system is validated by checking the answers to SPARQL queries.
In addition to the overall system, the system components were also evaluated. Inconsistencies in ontologies can be detected using ontological reasoners, which check whether there are contradictions in class hierarchies or class instances [4]. Reasoners such as HermiT and Pellet were used in many of the systems to evaluate the structural consistency of ontologies [79, 81, 84, 85, 86, 91, 92, 96, 107, 108, 75]. Additionally, some systems used ontology evaluation frameworks such as OntOlogy Pitfall Scanner! (OOPS!) [182], which was used by El-Sappagh et al. [84] and Zafeiropoulos et al. [75], and OQuaRE [183], which was used by Rhayem et al. [105]. Some systems also evaluated the effect of different components within the same system through ablation studies. For example, Ali et al. [77] tested the performance of their BiLSTM model for classifying healthcare data while using an ontology and without using an ontology. The results showed an increase in the accuracy of the model when combined with an ontology. Similarly, Ali et al. [78] compared the performance of their proposed ensemble deep learning model with and without feature selection. Additionally, systems that implemented ML used well-known metrics such as accuracy, precision, recall, F-score, and mean square error to evaluate the ML models.
6.4 Accessibility of research outputs
The sharing of research outputs, such as code, ontologies, knowledge graphs, and data, is a critical aspect of ensuring research is reproducible and verifiable. These resources can also be built upon by other researchers, contributing to their reuse for more efficient system development. This is severely neglected among the selected systems, with only four articles including publicly accessible links to their research outputs. Among them are Chatterjee et al. [81], who include their OWL ontology, simulated data, propositional variables, rule base, and queries as multimedia appendices. Similarly, Zafeiropoulos et al. [75] make their proposed ontology, queries, rules, code, and even research papers available via a GitHub repository. Using platforms like GitHub rather than static files has the advantage of version control, allowing researchers to manage future updates and revisions. De Brouwer et al. [83] include a link to a GitHub repository associated with the Data Analytics for Health and Connected Care ontology, which their proposed ontology extends and was developed by their research group. However, the new mBrain ontology reported in the article is not made available. In contrast, the ontology proposed by El-Sappagh et al. [84] has been published on Bioportal [^37] https://bioportal.bioontology.org, a popular repository of biomedical ontologies. However, no other research outputs, such as rules and queries, are made available.
Alti et al. [80] include a note that the data associated with their study is available by request. However, this is a suboptimal approach as it is impossible to guarantee the authors’ willingness or ability to consistently respond to such requests over time, potentially leading to prolonged delays or even a complete lack of response. Additionally, an important aspect of accessibility is ensuring that shared resources remain available in the long term. The article by Reda et al. [73] includes a link to a web portal developed as part of the system. According to the authors, the portal features a video tutorial and sample datasets for testing purposes. Unfortunately, as at the writing this article, the web portal is inaccessible and we are unable to ascertain if it was ever operational and for how long it may have been active. Finally, we note that it is understandable that researchers may be restricted in sharing participant data due to privacy concerns. A potential solution would be to seek participants’ consent in sharing their data anonymised and non-identifiable form. Table 12 indicates the links shared by researchers.
Table 12: Links to system outputs as shared by researchers.
System Type Link
Chatterjee et al. [81] Static files https://www.jmir.org/2021/4/e24656#app1
De Brouwer et al. [83] GitHub repository https://github.com/predict-idlab/DAHCC-Sources
El-Sappagh et al. [84] Bioportal https://bioportal.bioontology.org/ontologies/FASTO
Reda et al. [73] Web page - currently inaccessible http://137.204.74.19:8080/IFOPlatform/welcomePage.jsp
Zafeiropoulos et al. [75] GitHub repository https://github.com/KotisK/Wear4PDmove
6.5 Summary
This section has critically examined the quality of the selected systems, with a focus on four criteria: the data sources and devices used to collect the data; the development methodologies and tools used; the evaluation approaches and rigour; and the accessibility of research outputs. A summary list of the development tools development tools, evaluation approaches, and evaluation metrics used by the systems, can be found in Table 18 in the appendix. Though this analysis extends beyond Semantic Web technologies, we also consider several factors that are specific to Semantic Web technologies such as methodologies, languages, frameworks, and semantic databases. To summarise the section, we conduct an assessment of each system based on critical aspects related to the four quality criteria. We then discuss the results of the assessment, highlighting the aspects that are poorly addressed among the systems.
6.5.1 Quality assessment
Mirroring our assessment of how well the systems tackle the key challenges in Section 5.8, we have also evaluated the quality of the systems as reported in the corresponding research articles. We base our evaluation on the aspects of the quality criteria which are summarised in Table 13, and use the same five-point rating scale (✗, Low, Medium, High, and Very High) determined by the percentage of aspects that each system has met. The quality ratings for each system are shown in Table 19 in the appendix.
Table 13: Important aspects related to the quality evaluation criteria.
Criteria Aspects & scoring guide
Data and devices 1. Details are given regarding the source of sensor data and other health-related data used in developing or evaluating the systems, e.g.: whether existing or simulated datasets were used or new data collected, the nature of the data, the availability of the data, the number and description of the people the data was collected from.
0: no details on data sources are given; 1: only partial details are given; 2: comprehensive details are given.
2. Details of the specific sensor device(s) used to collect data are given. This means that at least one device is explicitly named if it is an existing or commercially available device, or described if it is a novel prototype. A description based on the type of data measured (e.g. temperature sensor or blood pressure monitor) is insufficient.
0: no specific device is mentioned; 1: at least one specific device is mentioned.
System and components development 1. A methodology has been followed for the development of Semantic Web technologies, other system components, and/or the system as a whole. The methodology can be existing or novel, but it must outline the steps followed in a systematic manner.
0: no methodology is mentioned; 1: a methodology is mentioned for at least one system component.
2. Details of the languages, platforms, tools, and other software used for the development of the system and its components are given.
0: no languages, platforms, tools, or other software is mentioned; 1: only partial details are given, e.g. for some system components but not others; 2: comprehensive details are given.
Rigour of evaluation 1. The individual system components (Semantic Web technologies or other techniques) are evaluated using appropriate methods e.g. the use of reasoners, competency questions, evaluation frameworks, and metrics like precision, recall or F1 score.
0: no evaluation of components is mentioned. 1: evaluation of one component is done and results are given. 2: more than one component is evaluated, and/or an established evaluation framework is used. 3: in addition to the evaluation of the individual components, the impact of the different components and the way the work together is evaluated, e.g. through ablation studies.
2. The system is compared with other approaches or systems (i.e. the state of the art) in a systematic way using formally defined criteria, and/or evaluation of the system is done by domain experts.
0: not done; 1: done
3. The potential real-world functionality of the system is evaluated, either through a use case scenario, case study, simulation, or deployment.
0: potential real world functionality is not considered; 1: a use case scenario/case study/simulation is used; 2: the system is deployed and/or user studies are carried out.
4. Non-functional requirements (NFRs) have been considered and details of this have been provided (e.g. scalability, adaptability, usability, security).
0: no NFRs have been considered or insufficient details are provided; 1: at least one NFR has been considered and explained.
Accessibility of system outputs Research outputs, including but not limited to code, ontologies, knowledge graphs, rules, queries, and data, have been made publicly and readily available without the need to contact the authors.
0: no resources are available; 1: at least one resource is readily available; 2: more than one resource is readily available.
6.5.2 Discussion
Table 14 shows the number of systems with a particular rating for each criteria, while the combined radar chart in Figure 4 provides a visualisation of the overall quality of the systems. Separate radar charts for the individual systems are also available [^38] https://public.flourish.studio/visualisation/17845323. The accessibility of research outputs is by far the most overlooked quality criterion, with only four systems making their resources publicly accessible. With regard to the rigour of evaluation, most systems achieved either a low or medium rating. This can be primarily attributed to the fact that numerous researchers only reported the evaluation of one system component, failing to evaluate or account for the impact of other components. The potential real-world functionality of the system was another poorly addressed aspect of evaluation. Most systems were not tested with actual users; rather, case studies and scenarios were the most common approach. Additionally, most researchers overlooked the importance of evaluating their systems against external benchmarks, such as drawing comparisons with similar existing systems or seeking evaluations from domain experts.
The description of data collection methods or existing datasets was generally well done among the systems, with many of them giving adequate details regarding the number and demographics of participants or dataset records and properly citing reused datasets. However, less than half of the systems gave details of the specific devices used to collect the sensor data. A likely explanation for this is that many of the proposed systems are theoretical proposals rather than functional implementations, and therefore they were not tested on real sensor data collected from actual devices. Finally, with regard to system development, most researchers adequately reported on the tools used to develop the different components of their proposed systems. However, only a small fraction of the researchers reported the use of an existing development methodology, or else adequately described the systematic steps taken to develop each system component.
Table 14: Counts of number of systems with each rating across the four main quality criteria.
Rating Criteria Data & devices System & components development Rigour of evaluation Accessibility of research outputs
11 1 2 39
Low 13 8 19 N/A
Medium 12 24 16 2
High N/A N/A 6 N/A
Very High 7 10 0 2
Figure 5: Combined radar chart showing the quality of the systems based on the specified criteria.
7 System architectures
The architecture of a system can be defined as an abstraction of the system in the form of a set of software structures needed to reason about it [184]. An important concept when discussing system architectures is the architectural style, which defines constraints on the form and structure of an architecture [185]. This is closely related to the architectural pattern, which is a reusable, well-established architectural solution to a recurring design problem [184]. As summarised in Table 4, the systems implement a range of architectural styles and patterns. This section will discuss the architectures of the systems, including how they support the achievement of the seven key challenges discussed in Section 5.
7.1 Architectural styles and patterns
7.1.1 Layered architecture
The most common type of architecture among the systems is the layered architecture, implemented in 24 of the systems [76, 77, 78, 79, 80, 85, 86, 87, 89, 90, 97, 98, 99, 100, 101, 102, 73, 106, 107, 108, 109, 72, 110, 112]. It is also the most common architectural pattern used in software systems generally [186, 187] and among sensor-based and IoT systems [45, 56]. In this pattern, each layer consists of a group of subtasks, with each group being at a particular level of abstraction [186]. This offers several advantages. It is simple to understand, and the separation of concerns among the different layers makes it easy to test and maintain the systems developed using this architecture [187]. Among the systems, there are variations in the number of layers and their functionality. However, the first layer is typically dedicated to data collection from wearable or ambient sensors as well as other data sources. It may be named the data collection layer, as in the systems by Ali et al. [77, 78], the sensing layer, as in the systems by Elhadj et al. [85] and Esposito et al. [86], or the user layer as in the system by Alti et al. [80]. Other typical layers include a data storage layer in which data is securely stored; networking layer which manages data communication and transmission in the system; inference and data analysis layer, in which the raw data is processed and analysed to derive important insights; and finally, presentation layer in the form of a user interface where individuals and in some cases, their clinicians and caregivers, can receive visualisations and alerts. Other specialised layers may also be included, such as the security layer in the system by Ali et al. [79], or the agents modelling and reasoning layer as proposed by Akhtar et al. [76].
7.1.2 Modular architecture
Similar to the layered architecture is the modular architecture, in which the system is subdivided into modules, blocks, or subsystems. This is the second most common architectural pattern among the systems, with some kind of modular pattern implemented in 19 of the systems [81, 82, 83, 84, 91, 93, 94, 95, 96, 99, 100, 103, 105, 74, 70, 75, 111, 112, 71]. Modular and layered architectural patterns can be used concurrently. For example, in the system proposed by Zhang et al. [112], the client management module has a middleware with a layered architecture. Additionally, because layered architectures tend to be monolithic, making them less agile and difficult to scale and deploy [187], modularity of layered architectures is advised, in which each layer consists of a modular set of components with a single function or purpose [188]. This is implemented by Mavropolous et al. [100], whose architecture has 3 levels (layers), with each containing specific modules. For example, the sensors management level contains a data analysis module, while the communication understanding level contains a natural language processing module. Similarly, Lopes de Souza et al. [99] implement a semantic module within their layered architecture.
7.1.3 Service-oriented architecture
Another well-known architectural pattern is the service-oriented architecture, a distributed pattern in which system components provide and consume services [184]. In service-oriented architectures, the different aspects of the challenges can be achieved using specialised services. For example, Hristoskova et al. [92] implement services such as a notification service to generate alerts (decision support) and a user location service to localize specific users (context awareness). While the service-oriented architectural pattern is powerful and offers a high level of abstraction, it is often overly complex and difficult to understand [187]. A way of mitigating these issues is to implement services in a layered architecture, as is done in several other systems [80, 87, 102, 72]. Additionally, agents can be used to effectively manage services, as is the case in the systems proposed by Alti et al. [80] and Fenza et al. [87].
7.1.4 Agent-based architecture
[…]
7.2 Proposed reference architecture
Based on an analysis of the systems as well as an overview of general sensor-based systems, a reference architecture for personal health monitoring systems is presented in Figure 6. The architecture consists of three layers as described below:
  1. 1.
    The data layer contains two modules. The data acquisition module supports the acquisition of data from body sensors as well as ambient sensors, health records, and user-submitted sources such as questionnaires and social media content. The data preprocessing module supports the preprocessing of the acquired data, including data cleaning, normalisation, and feature extraction.
  2. 2.
    The analysis and decision layer consists of the situation analysis module, which provides functionality to derive relevant detected and predicted situations from the data using techniques such as rules, ML, Bayesian networks, and fuzzy logic; and the decision support module, which follows up on the detected and predicted situations to recommend interventions that mitigate adverse situations and promote favourable ones. Central to both modules is expert health knowledge, including established clinical guidelines.
  3. 3.
    The presentation layer provides functionality through which users can receive communication from and interact with the system. The user communication module provides support for the system-generated communication of situations and recommended interventions through mediums such as text messages and emails, while the user interface module provides web and mobile applications with which users can interact with the system.
The architecture also includes a cross-cutting knowledge graph which represents heterogeneous health data. The underlying data schema is defined through an ontology, allowing for the semantic annotation of data and reuse of existing semantic resources such as the SAREF core ontology and its extensions for the health domain. The knowledge graph can be implemented and stored in graph databases such as GraphDB. A linked data approach is recommended to ensure related data is interlinked, thereby enabling data integration and reuse. This architecture is not only consistent with the layered architectures proposed in related reviews [49, 48, 46, 59, 56], but also includes modules to separate related but distinct functionalities within each layer, thereby mitigating the monolithicity of the layered approach. Table 15 highlights the key functionalities, recommended tools and techniques, and the key challenges addressed in each layer and module.
Figure 6: Reference architecture for sensor-based personal health monitoring systems.
Table 15: Summary of the layers and modules within the proposed reference architecture
Layer Module Inputs, processing tasks, and outputs Suggested development tools Challenges addressed
Presentation layer User communication module - Inputs: detected and predicted situations, recommended interventions, and explanations - Processing: content structuring and adaptation - Outputs: text messages and emails - Communication protocols e.g. Internet Message Access Protocol (IMAP), Simple Mail Transfer Protocol (SMTP), Short Message Service (SMS) - Messaging communication software e.g. Twilio, Plivo - Asynchronous message queuing software e.g. RabbitMQ, Apache Kafka - Templating engines - Interoperability - Context awareness - Explainability
User interface module - Inputs: software requirements, multimedia content, dynamic situations, interventions and explanations - Processing: development of software applic- ations in accordance with software development methodologies - Outputs: web and mobile interfaces and documentation - Web development languages e.g. HTML, CSS, Python, Javascript - Web development frameworks e.g. Python’s Django and Flask and Javascript’s Angular and React - Mobile development languages e.g. Kotlin, Swift, Java - Interoperability - Context awareness - Explainability
Analysis and decision layer Situation analysis module - Inputs: preprocessed health data, expert knowledge - Processing: execution of situation analysis rules, algorithms and models - Outputs: detected and predicted situations and explanations - Rule expression languages and extensions e.g. SHACL, SWRL, Bayes-SWRL - Semantic query languages e.g. SPARQL - Programming languages for ML model development and/or fuzzy logic implementation e.g. Python and MATLAB - Semantic Web editors and frameworks e.g. Protege and Apache Jena - Reasoners e.g. Pellet and HermiT - ML and deep learning libraries and frameworks e.g. Python’s scikit-learn, PyTorch, TensorFlow - Fuzzy logic development libraries e.g. MATLAB’s Fuzzy Logic Toolbox and Python’s Scikit-Fuzzy - Bayesian network modelling software e.g. Netica - Interoperability - Context awareness - Situation detection - Situation prediction - Explainability - Uncertainty handling
Decision support module - Inputs: detected and predicted situations, expert knowledge - Processing: execution of decision support rules, algorithms, and models - Outputs: recommended interventions and explanations - Interoperability - Context awareness - Decision support - Explainability - Uncertainty handling
Data layer Data acquisition module - Inputs: sensor devices, electronic health records, and user-submitted sources e.g. questionnaires, social media content - Processing: acquisition of data - Outputs: acquired heterogeneous health data - Data formatting languages e.g. XML, JSON, Turtle - Database query languages e.g. SQL - Data transmission protocols e.g. Bluetooth, Bluetooth Low Energy, ANT+ - Social media APIs - Interoperability - Context awareness
Data preprocessing module - Inputs: raw health data - Processing: data cleaning (handling missing, noisy, or erroneous data), signal processing, feature engineering - Outputs: preprocessed health data - Programming languages for data analysis and signal processing e.g. Python, MATLAB, R - Data analysis libraries e.g. Python’s NumPy, SciPy, Pandas - Data visualisation platforms e.g. Tableau, Microsoft Power BI Uncertainty handling
8 Discussion
8.1 Summary of findings
Figure 7 shows a map outlining the current state of the field. In the remainder of this section, the inadequacies and limitations in current systems will be highlighted, paving the way for opportunities for future research.
Figure 7: Map showing the current state of the field.
8.1.1 Summary of the extent to which key challenges are addressed
Our findings show that three of the seven key challenges are particularly poorly addressed among the systems: situation prediction, explainability, and uncertainty handling. Most of the systems included in this study do not adequately address the challenge of situation prediction, as they are incapable of predicting health risks or giving insight into how detected conditions may progress with time. In order to achieve the vision of precision health, it is important for health monitoring systems to go beyond detecting current health states and move towards the anticipation and mitigation of adverse health states. With regard to explainability, while all the reviewed systems use Semantic Web technologies, which are inherently interpretable, many of them do not present explicit explanations for system outputs. Additionally, none of the systems implement the criteria for good explanations as outlined by Molnar et al. [140], nor do any systems mention tailoring the explanations presented to suit different audiences. Uncertainty handling is similarly poorly implemented or not addressed at all in majority of the systems. While some of the systems consider the impact of sensor limitations such as noise and missing values, most do not address the inherent uncertainty present in situation analysis and decision support in the health domain. This hinders their ability to perform reliably when faced with ambiguous data or vague or limited knowledge, thus reducing their trustworthiness and dependability. Both situation prediction and uncertainty handling can be enhanced by a combination of techniques, as suggested by Behera et al. [65], such as ML and Bayesian networks. Few of the systems take such an approach, with the majority using solely rule-based reasoning.
To a lesser extent, interoperability and decision support are also not fully addressed in the selected systems. Considering interoperability, we found that only about a third of the systems take advantage of established sensor ontologies such as SAREF. Neglecting to use such ontologies limits the standardisation and expressiveness of the descriptions of sensors and, importantly, sensor data. This results in less effective querying of and reasoning on sensor data, which in turn negatively impacts situation analysis. In addition, while some systems incorporate established medical terminologies such as SNOMED CT and ICD, nearly all the systems fail to consider existing health data standards, such as FHIR. This can be attributed to the fact that there is a gap in tooling support for Semantic Web representations such as RDF with standards such as FHIR [189]. However, the lack of these integrations limits the extent to which such systems can use existing health data such as medical records. There is also significant room for improvement in addressing the challenge of decision support. While most of the included systems incorporate alerts to warn of hazardous situations, many do not offer recommendations or reminders for medication or lifestyle factors such as diet and exercise. Similarly, only a few of the systems report using established medical guidelines, which can help to provide a sound justification for any recommendations made, thereby enhancing the trustworthiness of the systems. However, the most overlooked aspect of decision support remains the human-centered aspect. Systems should support users’ agency to cognitively engage in decision-making by presenting them with various options and their potential outcomes, and allowing them to be the final decision-makers. While some of the selected systems do suggest recommendations, none offer more than one potential option or present their potential outcomes.
8.1.2 Summary of the quality assessment
With regard to the quality assessment, we found that 19 of the systems did not report the data collection methods or sources. Additionally, only 16 of the systems reported the specific devices used for data collection. It can be assumed that such systems may not have been properly validated using realistic data, which casts doubts on the claims made regarding the system functionality and performance. To mitigate this, researchers should clearly indicate which data was used to validate their systems, including how the data was collected, who it was collected from, and the devices that were used. Concerning the development of the systems, only 13 systems used an existing methodology or else systematically outlined the development steps followed for any of the system components. However, nearly all the systems provided details of the languages, platforms, tools, and other software used in the development process.
When it comes to the evaluation of the systems and components, nearly all the systems reported on the methods, metrics, and results of the evaluation process. However, only 15 included some kind of external evaluation, whether through a systematical comparison with other similar systems or seeking evaluation from domain experts. Furthermore, as was found in the review by Haque et al. [9], most of the selected systems are yet to be evaluated in real-world settings. While this is to be expected in an emerging area, it is imperative that more systems be evaluated in real-world settings going forward, so that practical challenges and user feedback can be identified early on and considered in future system proposals. This feedback loop is essential for undertaking further research into personal health monitoring systems that fully harness the potential of Semantic Web technologies. Finally, accessibility of resources was poor among the systems, with only four systems providing access to relevant system files. Wherever possible, researchers should include the research outputs such as ontologies, data, and code as publicly accessible supplementary material in order to enhance reproducibility and verifiability.
8.2 Future research directions
This study highlights the fact that many personal health monitoring systems do not fully leverage reusable resources, and instead opt to build resources from scratch. We also find that an overwhelming majority of the systems are not build with reusability in mind, as evidenced by the limited availability of research outputs from various researchers. Although different health conditions may require specific features and functionalities, Semantic Web technologies have the potential to be extendable, allowing for the addition of knowledge as it evolves and making them suitable for reuse across a wide range of health monitoring applications. We therefore invite researchers to not only reuse existing resources but also to build generalisable semantic and non-semantic system components and make them publicly available. This would play an important role in accelerating the development of personal health monitoring systems by avoiding the redundant efforts.
Another takeaway that has clearly emerged from this study is the advantage of combining Semantic Web technologies with other AI techniques such as ML and Bayesian networks. Integrating these approaches can significantly improve the tackling of the seven key challenges identified in the study. ML can also be leveraged to support the development of Semantic Web technologies [190, 191]. Furthermore, we encourage researchers to explore recent software libraries such as DeepOnto [192], which support ontology engineering tasks using deep learning tools and pre-trained language models. We also note that most systems do not report the use of recently-proposed Semantic Web standards, with only one system using SHACL and none using RDF-star and Notation 3. We encourage researchers to explore these and other state-of-the-art standards.
Additionally, a number of the systems do not take into account factors such as diet, exercise, and other determinants of health. The next generation of personal health monitoring systems must be more holistic, focusing not only on disease but also on overall wellness. This includes the monitoring of emotional and mental states, which has been shown to be linked to physical health [193]. Such information can be represented using Semantic Web technologies, including ontologies [194] and knowledge graphs [195]. The inadequately addressed challenges, together with the need for more holistic health monitoring, present interesting and important directions for future research in the field.
8.3 Limitations of this study
While we believe that this article offers a comprehensive overview of the use of Semantic Web technologies in personal health monitoring systems, it is a very broad area and thus we have necessarily had to delimit the scope of the article. We focus the depth of coverage on the seven key challenges, the quality analysis of the selected systems, and the proposed reference architecture. While a few additional challenges are discussed, they are not included in the in-depth analysis, and other potential challenges may not be mentioned in this article. Furthermore, although we provide an overview of sensors used for health monitoring and highlight some hardware-based interoperability challenges, the practical aspects of the seamless integration of physical sensors and the real-time processing of sensor data are not discussed in depth.
9 Conclusion
This systematic mapping study has analysed the landscape of sensor-based personal health monitoring systems that incorporate Semantic Web technologies. After a careful consideration of the pertinent issues in this application area, we identified seven key challenges that such systems must address. In a systematic process, we selected 43 systems as representative of the state of the art in the field, and critically analysed them based on their capacity to address the seven challenges. We also evaluated the quality of the research undertaken to develop them. Moreover, we discussed the architectures of the selected systems, and proposed a reference architecture to streamline the development of such systems. Lastly, we discussed the key findings of the study and highlighted opportunities for future research. It is our hope that this study will serve as a comprehensive overview of the field and spur further high-quality research in effective personal health monitoring systems.
{ack}
This work was financially supported by the Hasso Plattner Institute for Digital Engineering through the HPI Research School at the University of Cape Town. It was also supported in part by the National Research Foundation of South Africa (grant numbers: 151217, SRUG2204264808).
Appendix A Reused semantic resources
Table 16: Reused semantic resources, the systems that reuse them, and the challenges they address.
Reused resource Description Systems that reuse it Challenge(s) that it addresses
Amigo Device Ontology Provides support for the description of devices, user context, quality of service parameters, and communication protocols. [92] Interoperability; Context awareness
Association Ontology [^39] http://purl.org/ontology/ao/core Provides basic concepts and properties for describing specific association statements to something. [110] Interoperability
Basic Formal Ontology [^40] https://basic-formal-ontology.org A foundational ontology designed for use in supporting information retrieval, analysis, and integration in scientific and other domains. [84] Interoperability
Basic Geo (WGS84 - lat/long) [^41] https://www.w3.org/2003/01/geo A basic RDF vocabulary that provides a namespace for representing lat(itude), long(itude), and other information about spatially-located things. [73] Interoperability; Context awareness
BioMedBridges Diabetes Ontology Represents expert knowledge about stages and phenotypes of type 2 diabetes. [77] Interoperability; Situation detection; Situation prediction; Explainability
Chronic Obstructive Pulmonary Disease Ontology (COPDology) Contains concepts related to the disease, environment, equipment, patient data, and treatment. [100] Interoperability; Context awareness; Situation detection; Situation prediction; Decision support; Explainability
Cyc Ontology [^42] https://cyc.com/platform/ A large knowledge base of common sense and background knowledge. [104] Interoperability
Data Analytics for Health and Connected Care (DAHCC) Ontology Captures care, patient, daily life activity recognition and lifestyle domain knowledge. [83, 75] Interoperability; Context awareness; Situation detection; Situation prediction; Explainability
DBPedia A knowledge base of extracted structured information from Wikipedia in the form of an open knowledge graph served as Linked Data. [73, 72] Interoperability
Diabetes Mellitus Diagnosis Ontology A diabetes knowledge base that supports automatic reasoning for solving problems related to diabetes diagnosis. [84] Interoperability; Situation detection; Situation prediction; Decision support; Explainability
Diabetes Mellitus Treatment Ontology Provides knowledge about type 2 diabetes and its patients including complications, symptoms, tests, interactions, and treatment plans. [77, 84] Interoperability; Situation detection; Situation prediction; Decision support; Explainability
Diseasome A.k.a. the human disease network; a graph that captures all genetic disorders and disease genes and the links between them. [72] Interoperability; Situation detection; Situation prediction; Explainability
DOLCE+DnS Ultra Lite (DUL) Ontology A foundational ontology that provides a set of upper level concepts for easier interoperability among middle and lower level ontologies. [83, 74] Interoperability
DrugBank A knowledge resource for drug, drug–target and related pharmaceutical information. [72] Interoperability; Decision support; Explainability
Drug Target Ontology Provides formalized and standardized classifications and annotations of druggable protein targets. [77] Interoperability; Decision support; Explainability
Event Ontology Deals with the notion of reified events, which may have a location, a time, active agents, factors and products. [110] Interoperability; Context awareness
Fast Healthcare Interoperability Resources (FHIR) OWL Ontology OWL ontology for FHIR Resources represented in RDF. [84] Interoperability
FHIR & SSN-based T1 Diabetes Ontology (FASTO) Ontology integrating FHIR standard and SSN ontology with clinical practice guidelines for real time management of insulin for diabetes patients. [77] Interoperability; Situation detection; Situation prediction; Decision support; Explainability
Food Ontology A lightweight ontology for describing recipes, ingredients, menus, and diets. [106] Interoperability; Decision support; Explainability
Friend of a Friend (FOAF) Ontology An ontology describing persons, their activities, and their relations to other people and objects. [85, 90, 100, 73, 106, 107, 110] Interoperability; Context awareness
Table continued on next page.
Reused semantic resources, the systems that reuse them, and the challenges they address.
Reused resource Description Systems that reuse it Challenge(s) that it addresses
General User Model Ontology A user model ontology for use in user-adaptive or ubiquitous computing systems. [100] Interoperability; Context awareness
GeoNames Ontology Provides elements of description for geographical features. [105] Interoperability; Context awareness
HealthIoT Ontology Provides semantic representation of both the medical connected objects (i.e. sensors) and their data. [94] Interoperability; Context awareness
Heart Failure Ontology Defines heart-failure-relevant information including the causes and risk factors, signs and symptoms, diagnostic tests and results, and treatment. [92] Interoperability; Situation detection; Situation prediction; Decision support; Explainability
Human Disease Ontology Represents common and rare disease concepts. [77] Interoperability
Informed Consent Ontology Represents processes and information pertaining to obtaining informed consent in medical investigations. [110] Interoperability; Context awareness
International Classification for Nursing Practice (ICNP) Provides a standardised terminology that can be used to record the observations and interventions of nurses. [85, 90] Interoperability
International Classification of Diseases (ICD) A tool for recording, reporting and grouping conditions and factors that influence health. [106, 107, 70] Interoperability; Situation detection; Explainability
International Classification of Functioning, Disability and Health (ICF) A framework for describing and organising information on functioning and disability. [106] Interoperability; Situation detection; Explainability
International Classification of Headache Disorders (ICHD-3) A detailed hierarchical classification of all headache-related disorders. [83] Interoperability; Situation detection; Explainability
IoT-lite Ontology A lightweight ontology to represent IoT resources, entities, and services. [105] Interoperability; Context awareness
Logical Observation Identifiers Names and Codes (LOINC) An international standard for identifying health measurements, observations, and documents. [84] Interoperability
Medical Subject Headings (MeSH) A vocabulary for indexing, cataloging, and searching biomedical and health-related information. [73] Interoperability
MIMU-Wear Ontology Describes wearable sensor platforms consisting of mainstream magnetic and inertial measurement units. [94] Interoperability; Context awareness
Mobile Object Ontology Models temporal, spatial, and domain related information about mobile objects. [105] Interoperability; Context awareness
Medical Web Lifestyle Aggregator Ontology Represents user data from the web. [110] Interoperability; Context awareness
Ontology for Nutritional Studies A comprehensive resource for the description of concepts in the broader human nutrition domain. [77] Interoperability; Decision support; Explainability
OpenThesaurus A multilingual web-based thesaurus. [104] Interoperability
OwlSpeak Ontology Represents static and dynamic concepts related to spoken dialogue. [100] Interoperability; Context awareness
OWL-S Ontology Describes Semantic Web services for automatic service discovery, invocation, composition, and interoperation. [87, 92] Interoperability; Context awareness; Situation detection; Situation prediction; Decision support
OWL-Time Ontology Describes the temporal properties of resources. [83, 100, 105, 107, 110] Interoperability; Context awareness
Parkinson and Movement Disorder Ontology Captures neurological findings, treatment plans, and instruments used to evaluate various traits of Parkinson’s Disorder. [75] Interoperability; Context awareness; Situation detection; Situation prediction; Decision support; Explainability
Physical Activity Concept Ontology Captures various concepts that are required to describe one’s physical activity. [94] Interoperability; Context awareness; Situation detection; Situation prediction; Decision support; Explainability
Places Ontology A lightweight ontology for describing places of geographic interest. [110] Interoperability; Context awareness
Table continued on next page.
Reused semantic resources, the systems that reuse them, and the challenges they address.
Reused resource Description Systems that reuse it Challenge(s) that it addresses
RxNorm Provides normalized names for clinical drugs and links to other drug vocabularies. [84] Interoperability
SAREF for eHealth and Ageing Well (SAREF4EHAW) Ontology An extension of the SAREF ontology for applications related to eHealth and ageing well. [83, 99] Interoperability; Context awareness
SAREF for Wearables (SAREF4WEAR) Ontology An extension of the SAREF ontology for applications related to wearables. [89] Interoperability; Context awareness
Semantic Sensor Network (SSN)/Sensor, Observation, Sample, and Actuator (SOSA) Ontology Describes sensors and their capabilities, measurement processes, observations, and deployments. [81, 83, 84, 85, 90, 94, 105, 74, 107, 75] Interoperability; Context awareness
Smart Applications REFerence (SAREF) Ontology An ontology to enable semantic interoperability for smart appliances. [75] Interoperability; Context awareness
Smart Body Area Network (SmartBAN) Ontology Describes data related to sensors in BANs. [84] Interoperability; Context awareness
SWRL Temporal Ontology Defines defines a set of builtins that can be used to represent temporal information in SWRL rules. [84] Interoperability; Context awareness
Systematized Nomenclature of Medicine Clinical Terms (SNOMED CT) A collection of medical terms providing codes, terms, synonyms and definitions used in clinical documentation and reporting. [81, 84, 98, 99, 73, 105, 107, 71] Interoperability
Translational Medicine Ontology [110] Interoperability; Context awareness
Unified Medical Language System (UMLS) Provides a mapping structure among different health and biomedical terminologies, classifications, coding standards, and vocabularies. [104, 71] Interoperability
Vital Sign Ontology Provides a controlled structured vocabulary for describing vital sign measurement data, the processes of measuring vital signs, and the relevant anatomical entities. [84, 94] Interoperability
WordNet A lexical English language database of semantic relations between words, linking them into semantic relations. [104, 73] Interoperability
Appendix B Challenges assessment
Table 17: Summary of the extent to which each system addresses the seven key challenges.
# System Year Interoperability (score out of 6) Cont. awareness (score out of 7) Sit. detection (score out of 3) Sit. prediction (score out of 3) Dec. support (score out of 6) Explainability (score out of 6) Unc. handling (score out of 6) Overall (cumulative score out of 34)
1 Akhtar et al. [76] 2022 Low (1) High (5) Medium (2) ✗ (0) Medium (3) Medium (3) Low (1) Medium (15)
2 Ali et al. [77] 2021 Medium (3) Medium (4) Very High (3) ✗ (0) Low (2) Low (2) Low (1) Medium (15)
3 Ali et al. [78] 2020 Low (2) Medium (3) Medium (2) Low (1) Medium (3) Low (1) Medium (2) Medium (14)
4 Ali et al. [79] 2018 Low (1) Medium (3) Very High (3) ✗ (0) Medium (3) Low (2) Medium (2) Medium (14)
5 Alti et al. [80] 2022 Low (1) High (5) Medium (2) Low (1) Low (1) Low (2) ✗ (0) Low (12)
6 Chatterjee et al. [81] 2021 Medium (4) High (5) Medium (2) ✗ (0) Medium (2) Medium (3) ✗ (0) Medium (16)
7 Chiang and Liang [82] 2015 Low (1) High (5) Medium (2) Low (1) Medium (4) Medium (3) Low (1) Medium (17)
8 De Brouwer et al. [83] 2022 Medium (4) Medium (6) Very High (3) Low (1) Medium (3) Low (2) Low (1) Medium (20)
Table continued on next page.
Summary of the extent to which each system addresses the seven key challenges.
# System Year Interoperability (score out of 6) Cont. awareness (score out of 7) Sit. detection (score out of 3) Sit. prediction (score out of 3) Dec. support (score out of 6) Explainability (score out of 6) Unc. handling (score out of 6) Overall (cumulative score out of 34)
9 El-Sappagh et al. [84] 2019 High (5) High (5) Low (1) ✗ (0) Medium (3) Low (2) ✗ (0) Medium (16)
10 Elhadj et al. [85] 2021 High (5) High (5) Medium (2) ✗ (0) Low (2) Medium (4) ✗ (0) Medium (18)
11 Esposito et al. [86] 2018 Low (1) Medium (4) Very High (3) ✗ (0) Low (2) Low (2) Medium (2) Medium (14)
12 Fenza et al. [87] 2012 Low (2) Low (2) Very High (3) Low (1) ✗ (0) Low (1) Medium (2) Low (11)
13 Garcia-Valverde et al. [88] 2014 ✗ (0) Medium (3) Low (1) ✗ (0) Medium (2) Low (2) Low (1) Low (9)
14 Hadjadj and Halimi [89] 2021 Medium (3) High (5) Low (1) ✗ (0) Medium (3) Low (2) ✗ (0) Medium (14)
15 Henaien et al. [90] 2020 Medium (4) Medium (4) Low (1) ✗ (0) Medium (3) Low (2) ✗ (0) Medium (14)
16 Hooda and Rani [91] 2020 Low (1) Low (2) Medium (2) ✗ (0) Low (2) Low (2) Low (1) Low (10)
17 Hristoskova et al. [92] 2014 Medium (3) High (6) Medium (2) Low (1) Medium (3) Low (2) Low (1) Medium (18)
18 Hussain and Park [93] 2021 Low (2) Medium (3) Medium (2) Low (2) Medium (2) ✗ (0) Low (1) Low (12)
[…]
21 Khozouie et al. [96] 2018 Low (2) High (5) Low (1) ✗ (0) Low (1) Low (2) ✗ (0) Low (11)
22 Kim et al. [97] 2014 Low (1) Medium (3) Very High (3) ✗ (0) Low (1) Low (1) ✗ (0) Low (9)
23 Kordestani et al. [98] 2021 Medium (3) Medium (4) Medium (2) ✗ (0) Medium (4) Low (2) Medium (2) Medium (17)
24 Lopes de Souza et al. [99] 2023 Medium (3) Medium (4) Very High (3) ✗ (0) Medium (4) Low (2) ✗ (0) Medium (16)
25 Mavropoulos et al. [100] 2021 Low (2) High (5) Medium (2) ✗ (0) Medium (4) Medium (4) ✗ (0) Medium (17)
26 Mcheick et al. [101] 2016 ✗ (0) Medium (3) Low (1) Low (1) Low (1) Low (2) Low (1) Low (9)
27 Mezghani et al. [102] 2015 Low (1) Medium (3) Low (1) ✗ (0) Low (1) Low (1) Low (1) Low (8)
28 Minutolo et al. [103] 2016 Low (1) Medium (3) Very High (3) ✗ (0) ✗ (0) Low (1) Low (1) Low (9)
29 Peral et al. [104] 2018 Medium (4) Medium (3) Low (1) Medium (2) Medium (3) Low (1) ✗ (0) Medium (14)
30 Reda et al. [73] 2022 Low (1) High (5) Medium (2) Low (1) ✗ (0) Low (1) Low (1) Low (11)
31 Rhayem et al. [105] 2021 Medium (4) Very High (7) Medium (2) Low (1) Medium (3) Medium (3) Low (1) Medium (21)
32 Spoladore et al. [106] 2021 Low (2) Medium (4) Medium (2) ✗ (0) Medium (4) Low (2) ✗ (0) Medium (14)
33 Stavropoulos et al. [74] 2021 Low (1) Medium (4) Medium (2) ✗ (0) Low (2) Low (2) ✗ (0) Low (11)
34 Titi et al. [107] 2019 Medium (4) High (6) Medium (2) ✗ (0) High (5) Low (2) Low (1) Medium (20)
35 Vadillo et al. [108] 2013 Low (2) High (5) Low (1) ✗ (0) Medium (4) Low (2) ✗ (0) Medium (14)
36 Villarreal et al. [109] 2014 Low (1) Medium (4) Medium (2) ✗ (0) Medium (3) Medium (4) ✗ (0) Medium (14)
36 Villarreal et al. [109] 2014 Low (1) Medium (4) Medium (2) ✗ (0) Medium (3) Medium (4) ✗ (0) Medium (14)
37 Xu et al. [72] 2017 Low (2) Medium (3) Medium (2) ✗ (0) Low (2) Low (2) ✗ (0) Low (11)
38 Yu et al. [70] 2022 Low (2) Medium (4) Low (1) ✗ (0) High (5) Medium (4) ✗ (0) Medium (16)
39 Yu et al. [110] 2017 Low (2) Medium (4) Medium (2) Low (1) Low (1) Low (2) ✗ (0) Low (12)
40 Zafeiropoulos et al. [75] 2024 Medium (3) High (5) Very High (3) Low (1) Low (3) Low (2) ✗ (0) Medium (17)
41 Zeshan et al. [111] 2023 ✗ (0) High (5) Low (2) ✗ (0) Low (1) Low (2) ✗ (0) Low (10)
42 Zhang et al. [112] 2014 Low (2) Low (2) Medium (2) ✗ (0) Low (1) Low (2) ✗ (0) Low (9)
43 Zhou et al. [71] 2022 Medium (4) Medium (4) Low (1) Low (1) Medium (3) Low (1) Low (1) Medium (15)
Appendix C Development tools, evaluation approaches, and evaluation metrics
Table 18: Summary of development tools, evaluation approaches, and evaluation metrics used by the systems.
# System Semantic Web languages, standards, reasoners, and development frameworks Other programming languages, libraries, and frameworks Database systems, data repositories, and data analysis tools Evaluation approaches Evaluation metrics
1 Akhtar et al. [76] Protégé NetLogo None mentioned Case study; System simulation None mentioned
2 Ali et al. [77] OWL; Protégé Java; WEKA Amazon S3; Apache Pig; Apache Hadoop ML model performance; Comparison with SOTA; Ablation study Accuracy; Precision; Recall; RMSE; MAE
3 Ali et al. [78] OWL; Protégé; SWRL Java; WEKA None mentioned ML model performance; Comparison with SOTA; Ablation study Accuracy; Precision; Recall; F-score; RMSE; MAE
4 Ali et al. [79] Apache Jena; OWL; Pellet; Protégé; SPARQL; SWRL Java None mentioned Comparison with SOTA; Expert evaluation of system; Ontology validation Accuracy; Precision; Recall; F-score
5 Alti et al. [80] Apache Jena; Protégé; SWRL JADE; Java; Java Expert System Shell MySQL Case study; Comparison with SOTA Execution time; Optimality; Application lifetime; No. of discovered services
6 Chatterjee et al. [81] Apache Jena; Apache Jena Fuseki; HermiT; OWL; Protégé; RDF; SPARQL; SWRL None mentioned None mentioned Ontology validation Ontology reasoning time
7 Chiang and Liang [82] Apache Jena; Protégé; RDF C#; Java; MATLAB MySQL Use case scenarios; System simulation None mentioned
8 De Brouwer et al. [83] Apache Jena; Protégé; RDF; SPARQL None mentioned MongoDB System queries; Use case scenarios None mentioned
9 El-Sappagh et al. [84] HermiT; OWL; Pellet; Protégé; SPARQL; SWRL SQL SQLite Ontology validation; Comparison with SOTA; Comparison with expert opinion; Use case scenarios Correctness; Completeness; Extensibility; Conciseness; Organisational fitness
10 Elhadj et al. [85] Apache Jena; Apache Jena Fuseki; OWL; Pellet; Protégé; SPARQL; SWRL Java MongoDB Use case scenarios None mentioned
11 Esposito et al. [86] OWL; Pellet; Protégé Java SQLite Case study; ALMA method; Ontology validation Modifiability
12 Fenza et al. [87] OWL; SPARQL JADE None mentioned Case study; Comparison with logic-based matching Precision; Recall
13 Garcia-Valverde et al. [88] Apache Jena None mentioned None mentioned Use case scenario Accuracy; Precision; F-score
14 Hadjadj and Halimi [89] Apache Jena; OWL; Protégé; RDF; SPARQL; SWRL None mentioned None mentioned Use case scenario; Comparison with expert opinion Similarity between system and expert opinion
15 Henaien et al. [90] Protégé; SWRL WEKA None mentioned None mentioned None mentioned
16 Hooda and Rani [91] Apache Jena; OWL; Pellet; Protégé; RDF; SPARQL; SWRL None mentioned None mentioned Ontology validation None mentioned
17 Hristoskova et al. [92] OWL; Pellet; SWRL None mentioned None mentioned Expert evaluation; User evaluation; Ontology validation Performance; Scalability
18 Hussain and Park [93] Protégé; RDF None mentioned Apache ActiveMQ; Apache Hadoop; Apache Spark; Elasticsearch; MariaDB Comparison with SOTA AUC; Accuracy; Precision; Recall; Neg. predictive value
[…]
Table continued on next page.
Summary of development tools, evaluation approaches, and evaluation metrics used by the systems.
# System Semantic Web languages, standards, reasoners, and development frameworks Other programming languages, libraries, and frameworks Database systems, data repositories, and data analysis tools Evaluation approaches Evaluation metrics
[…]
21 Khozouie et al. [96] OWL; Pellet; Protégé Java None mentioned Use case scenario; Ontology validation; Expert evaluation None mentioned
22 Kim et al. [97] Apache Jena; OWL; Protégé SQL None mentioned Ontology validation; User evaluation Precision; Recall; F-score; Likert score
23 Kordestani et al. [98] None mentioned Drools None mentioned Comparison between BN and ML diagnosis; Use case scenarios F-score
24 Lopes de Souza et al. [99] RDF GraphQL; Java; Javascript; Node.js; React SQLite Ontology evaluation None mentioned
25 Mavropoulos et al. [100] OWL; SPARQL None mentioned GraphDB Comparison with SOTA; End user evaluation Accuracy; Precision; Recall; F-score; Likert scale
26 Mcheick et al. [101] None mentioned Netica None mentioned Use case scenarios Adaptability
27 Mezghani et al. [102] Apache Jena Fuseki; RDF; SPARQL SQL None mentioned Use case scenario None mentioned
28 Minutolo et al. [103] OWL; SPARQL None mentioned None mentioned Case study None mentioned
29 Peral et al. [104] None mentioned None mentioned None mentioned Case study Similarity between actual and predicted values
30 Reda et al. [73] OWL; RDF; RML; SPARQL; SWRL None mentioned None mentioned System simulation Not mentioned
31 Rhayem et al. [105] Apache Jena; OWL; SPARQL; SWRL Drools None mentioned Comparison with SOTA; Ontology validation F-score; Precision; Recall; Response time; Ontology coverage
32 Spoladore et al. [106] OWL; Protégé; SPARQL; SWRL None mentioned Stardog Use case scenarios None mentioned
33 Stavropoulos et al. [74] OWL; SHACL; SPARQL Python; Django GraphDB Use case scenarios; Focus group with clinicians Scalability; Likert scale
34 Titi et al. [107] Apache Jena; Pellet; Protégé; RDF; SPARQL; SWRL Java; Java Server Faces; PrimeFaces MySQL System queries None mentioned
35 Vadillo et al. [108] Apache Jena; OWL; Pellet; Protégé JADE None mentioned Use case scenarios; Ontology validation Processing time
36 Villarreal et al. [109] None mentioned Java MySQL Case study; ALMA method; User evaluation Response time; Usability; Recommendation suitability
37 Xu et al. [72] SPARQL None mentioned None mentioned Case study; Comparison with SOTA None mentioned
38 Yu et al. [70] None mentioned Python; spaCy; NetworkX None mentioned ML model performance; Use case scenarios AUC; Chronic Care Model criteria
39 Yu et al. [110] Apache Jena; OWL; RDF; SPARQL; SWRL None mentioned Virtuoso System queries None mentioned
40 Zafeiropoulos et al. [75] OWL; Pellet; Protégé; RDF; SPARQL; SWRL Owlready2; Python; PyTorch; RDFlib None mentioned Use case scenario; Ontology evelaution; ML model performance MAE; MSE; R-Squared
41 Zeshan et al. [111] Apache Jena; OWL; Protégé; SWRL None mentioned None mentioned Use case scenario Precision; Recall; Response time
42 Zhang et al. [112] OWL; Protégé; RDF; SPARQL; SWRL None mentioned None mentioned Use case scenario None mentioned
43 Zhou et al. [71] None mentioned None mentioned None mentioned Use case scenario None mentioned
ALMA - Architecture-Level Modifiability Analysis; AUC - Area Under the Curve; BN - Bayesian Network; CCM - Chronic Care Model; JADE - Java Agent DEvelopment Framework; MAE - Mean Absolute Error; ML - Machine Learning; OWL - Web Ontology Language; RDF - Resource Description Framework; RML - RDF Mapping Language; (R)MSE - (Root) Mean Square Error; SHACL - Shapes Constraint Language; SOTA - State of the Art; SPARQL - SPARQL Protocol and RDF Query Language; SWRL - Semantic Web Rule Language; WEKA - Waikato Environment for Knowledge Analysis
Appendix D Quality assessment
Table 19: Summary of the quality evaluation of each system.
# System Year Data & devices (score out of 3) System & components development (score out of 3) Rigour of evaluation (score out of 7) Accessibility of research outputs (score out of 2) Overall (score out of 15)
1 Akhtar et al. [76] 2022 ✗ (0) Medium (2) Low (1) ✗ (0) Low (3)
2 Ali et al. [77] 2021 Medium (2) Medium (2) Medium (4) ✗ (0) Medium (8)
3 Ali et al. [78] 2020 Medium (2) Medium (2) Medium (4) ✗ (0) Medium (8)
4 Ali et al. [79] 2018 Low (1) Medium (2) High (6) ✗ (0) Medium (9)
5 Alti et al. [80] 2022 Low (1) Medium (2) Low (2) ✗ (0) Low (5)
6 Chatterjee et al. [81] 2021 Medium (2) Very High (3) Medium (3) Very High (2) Medium (10)
7 Chiang and Liang [82] 2015 Low (1) Very High (3) Low (1) ✗ (0) Low (5)
8 De Brouwer et al. [83] 2022 Very High (3) Medium (2) Low (2) Medium (1) Medium (8)
9 El-Sappagh et al. [84] 2019 ✗ (0) Very High (3) Medium (4) Medium (1) Medium (8)
10 Elhadj et al. [85] 2021 ✗ (0) Very High (3) Medium (3) ✗ (0) Low (6)
11 Esposito et al. [86] 2018 Very High (3) Medium (2) High (5) ✗ (0) Medium (10)
12 Fenza et al. [87] 2012 ✗ (0) Medium (2) Low (2) ✗ (0) Low (4)
13 Garcia-Valverde et al. [88] 2014 Medium (2) Low (1) Low (2) ✗ (0) Low (5)
14 Hadjadj and Halimi [89] 2021 Medium (2) Very High (3) Medium (3) ✗ (0) Medium (8)
15 Henaien et al. [90] 2020 Low (1) Medium (2) ✗ (0) ✗ (0) Low (3)
16 Hooda and Rani [91] 2020 Low (1) Medium (2) Low (1) ✗ (0) Low (4)
17 Hristoskova et al. [92] 2014 Very High (3) Medium (2) Medium (3) ✗ (0) Medium (8)
18 Hussain and Park [93] 2021 Very High (3) Medium (2) Medium (4) ✗ (0) Medium (9)
[…]
21 Khozouie et al. [96] 2018 ✗ (0) Medium (2) Low (2) ✗ (0) Low (4)
22 Kim et al. [97] 2014 Medium (2) Medium (2) Medium (4) ✗ (0) Medium (8)
23 Kordestani et al. [98] 2021 Medium (2) Low (1) Low (2) ✗ (0) Low (5)
24 Lopes de Souza et al. [99] 2023 Low (1) Medium (2) Low (1) ✗ (0) Low (4)
25 Mavropoulos et al. [100] 2021 Very High (3) Low (1) High (6) ✗ (0) Medium (10)
26 Mcheick et al. [101] 2016 Low (1) Low (1) Low (2) ✗ (0) Low (4)
27 Mezghani et al. [102] 2015 ✗ (0) Low (1) Low (2) ✗ (0) Low (3)
28 Minutolo et al. [103] 2016 ✗ (0) Medium (2) Medium (3) ✗ (0) Low (5)
29 Peral et al. [104] 2018 Medium (2) Low (1) Low (2) ✗ (0) Low (5)
30 Reda et al. [73] 2022 Low (1) Medium (2) Low (1) ✗ (0) Low (4)
31 Rhayem et al. [105] 2021 Medium (2) Medium (2) High (5) ✗ (0) Medium (9)
32 Spoladore et al. [106] 2021 Low (1) Medium (2) Low (2) ✗ (0) Low (5)
33 Stavropoulos et al. [74] 2021 Very High (3) Medium (2) Medium (3) ✗ (0) Medium (8)
34 Titi et al. [107] 2019 ✗ (0) Very High (3) Low (1) ✗ (0) Low (4)
35 Vadillo et al. [108] 2013 Low (1) Very High (3) Medium (3) ✗ (0) Medium (7)
36 Villarreal et al. [109] 2014 Medium (2) Low (1) High (6) ✗ (0) Medium (9)
37 Xu et al. [72] 2017 ✗ (0) Low (1) Medium (3) ✗ (0) Low (4)
38 Yu et al. [70] 2022 Medium (2) Medium (1) High (5) ✗ (0) Medium (8)
39 Yu et al. [110] 2017 Low (1) Very High (3) Low (1) ✗ (0) Low (5)
40 Zafeiropoulos et al. [75] 2024 Medium (2) Very High (3) Medium (3) Very High (2) Medium (10)
41 Zeshan et al. [111] 2023 Low (1) Medium (2) Medium (3) ✗ (0) Low (6)
42 Zhang et al. [112] 2014 Low (1) Medium (2) Low (2) ✗ (0) Low (5)
43 Zhou et al. [71] 2022 ✗ (0) ✗ (0) Low (1) ✗ (0) Low (1)